### v4

# Spectral bounds for certain special type of rational matrices

Abstract.
We derive bounds on the moduli of the eigenvalues of special type of rational matrices using the following techniques/methods: (1) the Bauer-Fike theorem on an associated block matrix of the given rational matrix, (2) by associating a real rational function, along with Rouché theorem for the rational matrix and (3) by a numerical radius inequality for a block matrix for the rational matrix. These bounds are compared when the coefficients are unitary matrices. Numerical examples are given to illustrate the results obtained.
Key words and phrases:
Rational matrices; rational eigenvalue problems; spectral bounds for rational matrices; matrix-valued functions; matrix polynomials; numerical radius.
2010 Mathematics Subject Classification:
15A18, 15A42, 15A22, 93B18, 47A56
1. Introduction
Rational eigenvalue problems, abbreviated henceforth as REPs, are an important class of nonlinear eigenvalue problems that arise in applications to science and engineering. REPs arise for instance, in applications to computing damped vibration modes of an acoustic fluid confined in a cavity [7], describing the eigenvibration of a string with a load of mass attached by an elastic string [4] and in application to photonic crystals [9]. Recent studies on REP can also be found in [14], [17], [8], and [16] and in the references cited in these. There are several spectral bounds for matrix polynomials depending on norms of coefficients and roots of associated scalar polynomials (see for instance, [11] and [6]). However, among the ones available in the literature, there is no easy or best way to determine the location of eigenvalues of rational matrices. For instance, one approach is to convert the rational matrix to a matrix polynomial and use existing results. However, in practice, it is difficult to determine the coefficients of the matrix polynomial that was obtained.
The purpose of this work is to derive bounds on the eigenvalues of rational matrices of special type. We provide bounds that can be calculated with a small computational effort. We work either over the field $\mathbb{C}$ of complex numbers or over the field $\mathbb{R}$ of real numbers. The vector space of $n\times n$ matrices over $\mathbb{C}$ (respectively, $\mathbb{R}$) is denoted by $M_{n}(\mathbb{C})$ (respectively, $M_{n}(\mathbb{R})$). $||\cdot||_{2}$ denotes the spectral norm of a square matrix. The condition number of a square matrix $A$ is denoted by $\kappa(A)$ is defined as $\kappa(A)=\begin{cases}||A||||A^{-1}||&\text{if $A$ is invertible}\\ \infty&\text{otherwise}\end{cases}$, where |||||$\cdot$||| | ⋅ | | is any matrix norm.
The organization of the manuscript is as follows. Section 2.1 contains a brief introduction to rational matrices. Some preliminary notions about rational matrices are provided. A bound on the eigenvalues of rational matrices of special type using an associated block matrix (see Section 2.1 for the definition) is given in Section 2.2. In Sections 2.3 and 2.4, bounds on the eigenvalues are given using roots of real rational functions and polynomials, respectively. A bound using the numerical range is given in Section 2.5. These bounds are compared in Section 3. Numerical illustrations are given in Section 4. The computations are done using Matlab.
2. Main results
The main results are presented in this section. We start with preliminaries on rational matrices. In the subsections that follow, we derive various bounds on the eigenvalues of rational matrices.
2.1. Rational matrices.
An $n\times n$ rational matrix is a matrix whose entries are rational functions, which can be expressed in the form $T(\lambda)=P(\lambda)-\sum_{i=1}^{k}\frac{s_{i}(\lambda)}{q_{i}(\lambda)}E_{i}$, where $P(\lambda)$ is an $n\times n$ matrix polynomial of degree $d$, $s_{i}(\lambda)$ and $q_{i}(\lambda)$ are scalar polynomials of degree $n_{i}$ and $d_{i}$ respectively, and $E_{i}$’s $\in M_{n}(\mathbb{C})$. Rational matrices are often known as rational matrix functions or matrix rational functions in the literature. An $n\times n$ rational matrix $T(\lambda)$ is said to be regular if its determinant does not vanish identically. The rational eigenvalue problem (REP) is to find a scalar $\lambda_{0}\in\mathbb{C}$ and a nonzero vector $v\in\mathbb{C}^{n}$ such that $T(\lambda_{0})v=0$, with $T(\lambda)$ being regular and $T(\lambda_{0})$ bounded (that is, $T(\lambda_{0})$ has finite entries). The scalar $\lambda_{0}$ so obtained is called an eigenvalue of $T(\lambda)$ and the vector $v$ is called an eigenvector of $T(\lambda)$ corresponding to the eigenvalue $\lambda_{0}$. Note that if $T(\lambda)=B$, where $B$ is a nonsingular matrix, then no complex number is an eigenvalue for $T(\lambda)$.
The nonlinear eigenvalue problem (abbreviated as NEP) seeks to find a scalar $\lambda_{0}$ and a nonzero vector $v\in\mathbb{C}^{n}$ satisfying $G(\lambda_{0})v=0$, where $G(\lambda)$ is a regular matrix-valued function ($G(\lambda)$ is square and its determinant does not vanish identically), and each entry of $G(\lambda_{0})$ is bounded. The majority of the nonlinear eigenvalue problems in applications are of the form, $G(\lambda)=-B_{0}+A_{0}\lambda+A_{1}f_{1}(\lambda)+\dots+A_{p}f_{p}(\lambda)$, where $f_{i}:\Omega\rightarrow\mathbb{C}$ are analytic functions and $\Omega$ is a region in $\mathbb{C}$. In [16], to study the eigenvalue problem of $G(\lambda)$ authors consider the $surrogate$ problem, $T(\lambda)v=\Big{(}-B_{0}+A_{0}\lambda+\displaystyle\sum_{i=1}^{m}\frac{B_{i}}{\lambda-\alpha_{i}}\Big{)}v=0$, where $\alpha_{i}$’s are distinct complex numbers. This is achieved by approximating $f_{i}$’s by a rational function of the form $r_{i}(\lambda)=\displaystyle\sum_{i=1}^{m}\frac{\sigma_{ij}}{\lambda-\alpha_{i}}$. The $\alpha_{i}$’s are the same for all functions $r_{i}$’s. This motivates us to study the rational matrices of the form
$T(\lambda)=-B_{0}+A_{0}\lambda+\frac{B_{1}}{\lambda-\alpha_{1}}+\dots+\frac{B_{m}}{\lambda-\alpha_{m}},$ (2.1)
where $B_{i}$’s are $n\times n$ matrices and $\alpha_{i}$’s are distinct complex numbers. Since we are interested in finding bounds on the eigenvalues of $T(\lambda)$ we assume $A_{0}$ to be nonsingular. We assume $A_{0}=I$, the identity matrix as one can multiply Equation (2.1) by $A_{0}^{-1}$. We, thus, consider rational matrices of the form
$T(\lambda)=\displaystyle-B_{0}+I\lambda+\frac{B_{1}}{\lambda-\alpha_{1}}+\frac{B_{2}}{\lambda-\alpha_{2}}+\cdots+\frac{B_{m}}{\lambda-\alpha_{m}},$ (2.2)
where $\alpha_{i}$’s are distinct complex numbers ordered $|\alpha_{1}|<|\alpha_{2}|<\cdots<|\alpha_{m}|$ and the $B_{i}$’s are $n\times n$ complex matrices. It turns out that many rational eigenvalue problems can be converted to this form (see for instance, Example 4.1). Note that for $T(\lambda)$ given in Equation (2.2), the REP can be converted to a linear eigenvalue problem, $P(\lambda)v=0$, where $P(\lambda)=I\lambda-C_{T}$ with $C_{T}=\begin{bmatrix}\alpha_{1}I&0&\cdots&0&-I\\ 0&\alpha_{2}I&\cdots&0&-I\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&\alpha_{m}I&-I\\ B_{1}&B_{2}&\cdots&B_{m}&B_{0}&\end{bmatrix}$ of size $(m+1)n\times(m+1)n$. $P(\lambda)$ is also a polynomial system matrix of $T(\lambda)$ with the state matrix $A(\lambda)=\text{diag}\big{(}(\lambda-\alpha_{1})I,\ldots,(\lambda-\alpha_{m})I\big{)}$ (see [8, 3] for details). Interestingly, the linearization of $T(\lambda)$ given in [2] using Fiedler matrices is the same as $P(\lambda)$. Thus, corresponding to $T(\lambda)$ we associate the block matrix $C_{T}$. Note that the eigenvalues of $T(\lambda)$ are also eigenvalues of $C_{T}$. Moreover, if all the $B_{i}$’s are nonsingular, then $T(\lambda)$ and $C_{T}$ have the same eigenvalues (see [8] for details). It is easy to verify that if one of the coefficients $B_{i}$ is singular, then the corresponding pole $\alpha_{i}$ is an eigenvalue of $C_{T}$. Therefore, $T(\lambda)$ has at most $(m+1)n$ eigenvalues.
2.2. Bound on the eigenvalues of $T(\lambda)$ using Bauer-Fike theorem
One of the well known results in the perturbation theory of the eigenvalue of a diagonalizable matrix is due to Bauer and Fike (Theorem 6.3.2, [12]). We make use of this result to find a bound for the eigenvalues of a rational matrix given in Equation (2.2). We state the Bauer-Fike theorem below.
Theorem 2.1.
Let $A\in M_{n}(\mathbb{{C}})$ be diagonalizable, and suppose that $A=S\Lambda S^{-1}$, in which $S$ is nonsingular and $\Lambda$ is diagonal. Let $E\in M_{n}(\mathbb{C})$ and |||||$\cdot$||| | ⋅ | | be a matrix norm on $M_{n}(\mathbb{C})$ that is induced by an absolute norm on $\mathbb{C}^{n}$. If $\hat{\lambda}$ is an eigenvalue of $A+E$, then there is an eigenvalue $\lambda$ of $A$ such that $|\hat{\lambda}-\lambda|\leq\kappa(S)||E||$ in which $\kappa(\cdot)$ is the condition number with respect to the matrix norm.
We now derive eigenvalue bounds for rational matrices given in Equation (2.2) using the Bauer-Fike theorem.
Theorem 2.2.
Let $T(\lambda)$ be as in Equation (2.2), with an eigenvalue $\lambda_{0}$. Then $|\lambda_{0}|\leq||E||+|\alpha_{m}|$, where $E=\begin{bmatrix}0&\cdots&0&-I\\ \vdots&\ddots&\vdots&\vdots\\ 0&\cdots&0&-I\\ B_{1}&\cdots&B_{m}&B_{0}\\ \end{bmatrix}$ and |||||$\cdot$||| | ⋅ | | is any matrix norm induced by an absolute norm on $\mathbb{C}^{n}$.
Proof.
We know that the eigenvalues of $T(\lambda)$ are also the eigenvalues of the associated block matrix $C_{T}$. Let us write $C_{T}$ as a sum of two matrices $A$ and $E$, where $A=\begin{bmatrix}\alpha_{1}I&\cdots&0&0\\ \vdots&\ddots&\vdots&\vdots\\ 0&\cdots&\alpha_{m}I&0\\ 0&\cdots&0&0\end{bmatrix}$ and $E=\begin{bmatrix}0&\cdots&0&-I\\ \vdots&\ddots&\vdots&\vdots\\ 0&\cdots&0&-I\\ B_{1}&\cdots&B_{m}&B_{0}\end{bmatrix}$. Since $A$ is a diagonal matrix, Theorem 2.1 guarantees that there exists an eigenvalue $\mu$ of $A$ such that $|\lambda_{0}-\mu|\leq\kappa(S)||E||$. Note that the similarity matrix $S$ in this case, is $S=I$, the identity matrix so that $\kappa(S)=1$. Therefore, $|\lambda_{0}-\mu|\leq||E||$. This implies $|\lambda_{0}|\leq||E||+|\mu|$. Since the eigenvalues of $A$ are $0,\alpha_{1},\dots,\alpha_{m}$, we see that $|\lambda_{0}|\leq||E||+|\alpha_{m}|$. This completes the proof. ∎
In particular, if we assume the $B_{i}$’s to be unitary matrices and the induced norm to be the spectral norm, then we get a bound which depends only on the number of poles of $T(\lambda)$ and their moduli. We state this below.
Theorem 2.3.
Let $T(\lambda)=-B_{0}+I\lambda+\frac{B_{1}}{\lambda-\alpha_{1}}+\cdots+\frac{B_{m}}{\lambda-\alpha_{m}}$, where $B_{i}$’s are $n\times n$ unitary matrices and $\alpha_{i}$’s are distinct complex numbers. Then $|\lambda_{0}|\leq\displaystyle\Bigg{\{}\frac{(2m+1)+(4m+1)^{1/2}}{2}\Bigg{\}}^{1/2}+|\alpha_{m}|$ for any eigenvalue $\lambda_{0}$ of $T(\lambda)$.
Proof.
By Theorem 2.2, we have $|\lambda_{0}|\leq||E||_{2}+|\alpha_{m}|$. To find $||E||_{2}$, consider $EE^{\ast}=\begin{bmatrix}I&\cdots&I&-B_{0}^{*}\\ \vdots&\ddots&\vdots&\vdots\\ I&\cdots&I&-B_{0}^{*}\\ -B_{0}&\cdots&-B_{0}&(m+1)I\end{bmatrix}$. Consider the matrix $Q(\lambda)$, defined by $Q(\lambda):=\begin{bmatrix}I&0&\cdots&0&0\\ -(I-I\lambda)^{-1}&I&\cdots&0&0\\ -(2I-I\lambda)^{-1}&-(2I-I\lambda)^{-1}&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ -\big{(}(m-1)I-I\lambda\big{)}^{-1}&-\big{(}(m-1)I-I\lambda\big{)}^{-1}&\cdots&I&0\\ B_{0}(mI-I\lambda)^{-1}&B_{0}(mI-I\lambda)^{-1}&\cdots&B_{0}(mI-I\lambda)^{-1}&I\\ \end{bmatrix}$. Observe that $\text{det}Q(\lambda)=1$ and $Q(\lambda)(EE^{*}-I\lambda)$ is a block upper triangular matrix with diagonal blocks $I-I\lambda,I-I\lambda-(I-I\lambda)^{-1},\dots,I-I\lambda-(m-1)\big{(}(m-1)I-I\lambda\big{)}^{-1}$ and $(m+1)I-I\lambda-m(mI-I\lambda)^{-1}$ respectively. Therefore, $\text{det}(EE^{*}-I\lambda)$ equals $\text{det}(Q(\lambda)(EE^{*}-I\lambda))$, which in turn is the product of the determinant of these diagonal blocks. On simplifying, we get
$\text{det}(EE^{*}-I\lambda)$
$=(1-\lambda)^{n}\Big{(}(1-\lambda)-\frac{1}{1-\lambda}\Big{)}^{n}\cdots$
$\hskip 85.35826pt\cdots\Big{(}(1-\lambda)-\frac{(m-1)}{(m-1)-\lambda}\Big{)}^{n}\Big{(}(m+1-\lambda)-\frac{m}{m-\lambda}\Big{)}^{n}$
$=(1-\lambda)^{n}\frac{\lambda^{n}(\lambda-2)^{n}}{(1-\lambda)^{n}}\cdots\frac{\lambda^{n}(\lambda-m)^{n}}{(m-1-\lambda)^{n}}\frac{(\lambda^{2}-(2m+1)\lambda+m^{2})^{n}}{(m-\lambda)^{n}}$
$=(-1)^{(m-1)n}\lambda^{(m-1)n}\big{(}\lambda^{2}-(2m+1)\lambda+m^{2}\big{)}^{n}.$
Thus, the eigenvalues of $EE^{*}$ are 00 and $\frac{(2m+1)\pm(4m+1)^{1/2}}{2}$, from which we see that $||E||_{2}=\displaystyle\Bigg{\{}\frac{(2m+1)+(4m+1)^{1/2}}{2}\Bigg{\}}^{1/2}$. ∎
2.3. Bounds on the eigenvalues of $T(\lambda)$ using rational functions.
Another method for determining bounds on the eigenvalues is to use norms of the coefficient matrices of $T(\lambda)$ and define a rational function whose roots are bounds for the eigenvalues of $T(\lambda)$. We set out to do this in this section. We begin with the following elementary lemma, whose proof is a simple consequence of the intermediate value theorem.
Lemma 2.4.
Let $q(x)=x-a_{0}-\frac{a_{1}}{x-b_{1}}-\dots-\frac{a_{m}}{x-b_{m}}$ be a real rational function, where the $a_{i}$’s are positive and $b_{i}$’s are nonnegative real numbers such that $b_{1}<b_{2}<\dots<b_{m}$. Then $q(x)$ has roots $R_{1},R_{2},\dots,R_{m+1}$ such that $R_{1}<b_{1}<R_{2}<b_{2}<\dots<R_{m}<b_{m}<R_{m+1}$.
Notice that since $q(x)$ is a rational function it has exactly $m+1$ roots. Moreover, $R_{i}$’s are positive for all $12\leq i\leq m+$ 1; however, $R_{1}$ can be less than or equal to zero. Now we shall find an upper bound for the eigenvalues of a rational matrix.
Theorem 2.5.
Let $T(\lambda)$ be as given in Equation (2.2), with an eigenvalue $\lambda_{0}$. Define a real rational function associated with $T(\lambda)$ as follows:
$q(x)=x-||B_{0}||-\frac{||B_{1}||}{x-|\alpha_{1}|}-\dots-\frac{||B_{m}||}{x-|\alpha_{m}|},$ (2.3)
where |||||$\cdot$||| | ⋅ | | is any induced matrix norm. Then $|\lambda_{0}|\leq R$, where $R$ is a real positive root of $q(x)$ such that $|\alpha_{i}|<R$ for all $i=1,2,\cdots,m$.
Proof.
By Lemma 2.4, there is a positive real root $R$ of $q(x)$ such that $|\alpha_{i}|<R$ for all $i=1,2,\dots,m$. Therefore,
00 0 $=q(R)=R-||B_{0}||-\frac{||B_{1}||}{R-|\alpha_{1}|}-\dots-\frac{||B_{m}||}{R-|\alpha_{m}|}$
$\implies R=||B_{0}||+\frac{||B_{1}||}{R-|\alpha_{1}|}+\dots+\frac{||B_{m}||}{R-|\alpha_{m}|}$
$\implies 1=R^{-1}||B_{0}||+R^{-1}\frac{||B_{1}||}{R-|\alpha_{1}|}+\dots+R^{-1}\frac{||B_{m}||}{R-|\alpha_{m}|}.$
By contradiction, suppose $|\lambda_{0}|>R$. Then $|\lambda_{0}-\alpha_{i}|\geq|\lambda_{0}|-|\alpha_{i}|>R-|\alpha_{i}|$. Since $|\alpha_{i}|<R$, we have $\frac{1}{|\lambda_{0}-\alpha_{i}|}<\frac{1}{R-|\alpha_{i}|}$ for all $i=1,2,\cdots,m$. For a unit eigenvector $v\in\mathbb{C}^{n}$ corresponding to $\lambda_{0}$, we have,
$\Big{|}\Big{|}-\lambda_{0}^{-1}B_{0}v+\frac{\lambda_{0}^{-1}B_{1}v}{\lambda_{0}-\alpha_{1}}+\dots+\frac{\lambda_{0}^{-1}B_{m}v}{\lambda_{0}-\alpha_{m}}\Big{|}\Big{|}$
$\leq|\lambda_{0}|^{-1}||B_{0}||+\frac{|\lambda_{0}|^{-1}||B_{1}||}{|\lambda_{0}-\alpha_{1}|}+\dots+\frac{|\lambda_{0}|^{-1}||B_{m}||}{|\lambda_{0}-\alpha_{m}|}$
$<R^{-1}||B_{0}||+\frac{R^{-1}||B_{1}||}{R-|\alpha_{1}|}+\dots+\frac{R^{-1}||B_{m}||}{R-|\alpha_{m}|}$
$=1.$
Then,
$0=||T(\lambda_{0})v||$ $=\Big{|}\Big{|}-B_{0}v+I\lambda_{0}v+\frac{B_{1}v}{\lambda_{0}-\alpha_{1}}+\dots+\frac{B_{m}v}{\lambda_{0}-\alpha_{m}}\Big{|}\Big{|}$
$=|\lambda_{0}|\Big{|}\Big{|}v-\lambda_{0}^{-1}B_{0}v+\frac{\lambda_{0}^{-1}B_{1}v}{\lambda_{0}-\alpha_{1}}+\dots+\frac{\lambda_{0}^{-1}B_{m}v}{\lambda_{0}-\alpha_{m}}\Big{|}\Big{|}$
$\geq|\lambda_{0}|\Big{(}1-\Big{|}\Big{|}-\lambda_{0}^{-1}B_{0}v+\frac{\lambda_{0}^{-1}B_{1}v}{\lambda_{0}-\alpha_{1}}+\dots+\frac{\lambda_{0}^{-1}B_{m}v}{\lambda_{0}-\alpha_{m}}\Big{|}\Big{|}\Big{)}$
$>|\lambda_{0}|(1-1)=0,$
a contradiction. This completes the proof. ∎
We now associate another rational function very similar to $q(x)$ to obtain a lower bound for the eigenvalues of rational matrices with added assumptions. We make use of a Rouché type theorem for analytic matrix-valued functions (Theorem $2.3$ of [6]) stated below.
Theorem 2.6.
Let $A,B:G\rightarrow\mathbb{C}^{n\times n}$, where $G$ is an open connected subset of $\mathbb{C}$, be analytic matrix-valued functions. Assume that $A(\lambda)$ is invertible on the simple closed curve $\gamma\subseteq G$. If $||A(\lambda)^{-1}B(\lambda)||<1$ for all $\lambda\in\gamma$, then $A+B$ and $A$ have the same number of eigenvalues inside $\gamma$, counting multiplicities. The norm is the matrix norm induced by any norm on $\mathbb{C}^{n}$.
The following theorem gives a lower bound for the eigenvalues of rational matrices.
Theorem 2.7.
Let $T(\lambda)$ be as in Equation (2.2). If $B_{0}$ is invertible and $||B_{0}^{-1}||^{-1}>\displaystyle\frac{||B_{1}||}{|\alpha_{1}|}+\dots+\frac{||B_{m}||}{|\alpha_{m}|}$, then $\tilde{R}\leq|\lambda_{0}|$ for any eigenvalue $\lambda_{0}$ of $T(\lambda)$, where $\tilde{R}$ is the unique positive root of the real rational function $p(x)=x-||B_{0}^{-1}||^{-1}-\frac{||B_{1}||}{x-|\alpha_{1}|}-\dots-\frac{||B_{m}||}{x-|\alpha_{m}|}$ such that $\tilde{R}<|\alpha_{i}|$ for all $i=1,2,\dots,m$.
Proof.
From Lemma 2.4, $p(x)$ has a unique root $\tilde{R}$ such that $\tilde{R}<|\alpha_{1}|$. Note that $p(0)=-||B_{0}^{-1}||^{-1}+\frac{||B_{1}||}{|\alpha_{1}|}+\dots+\frac{||B_{m}||}{|\alpha_{m}|}<0$ by the given hypothesis. Therefore, by the intermediate value theorem $0<\tilde{R}<|\alpha_{1}|$. Let $A(\lambda):=-B_{0}$ and $B(\lambda):=\displaystyle I\lambda+\frac{B_{1}}{\lambda-\alpha_{1}}+\dots+\frac{B_{m}}{\lambda-\alpha_{m}}$. Then $T(\lambda)=A(\lambda)+B(\lambda)$. Taking $G$ to be the disk $D(0,\tilde{R}):=\{z\in\mathbb{C}:|z|<\tilde{R}\}$, we see that $A(\lambda)$ and $B(\lambda)$ are analytic matrix-valued functions on $G$. Since $p(0)<0$ and $p(\tilde{R})=0$, we have $p(x)<0$ for all $0\leq x<\tilde{R}$. Therefore, for all $|\lambda|<\tilde{R}$ we have
$|\lambda|-||B_{0}^{-1}||^{-1}-\frac{||B_{1}||}{|\lambda|-|\alpha_{1}|}-\dots-\frac{||B_{m}||}{|\lambda|-|\alpha_{m}|}<0.$ (2.4)
Now for $|\lambda|<\tilde{R}$, consider
$||B(\lambda)||$ $=\Big{|}\Big{|}I\lambda+\frac{B_{1}}{\lambda-\alpha_{1}}+\dots+\frac{B_{m}}{\lambda-\alpha_{m}}\Big{|}\Big{|}$
$\leq|\lambda|+\frac{||B_{1}||}{|\lambda-\alpha_{1}|}+\dots+\frac{||B_{m}||}{|\lambda-\alpha_{m}|}$
$\leq|\lambda|+\frac{||B_{1}||}{|\alpha_{1}|-|\lambda|}+\dots+\frac{||B_{m}||}{|\alpha_{m}|-|\lambda|}\hskip 28.45274pt(\text{since}\,|\lambda|<|\alpha_{i}|)$
$=|\lambda|-\frac{||B_{1}||}{|\lambda|-|\alpha_{1}|}-\dots-\frac{||B_{m}||}{|\lambda|-|\alpha_{m}|}$
[…]
For any $\epsilon>0$, define $\gamma:=(\tilde{R}-\epsilon)e^{i\theta}$, where $0\leq\theta\leq 2\pi$. Then $||A^{-1}(\lambda)B(\lambda)||<1$ for all $\lambda\in\gamma$. Since $\epsilon>0$ is arbitrary, we see from Theorem 2.6 that the number of eigenvalues of $A(\lambda)$ and $A(\lambda)+B(\lambda)$ are same inside $D(0,\tilde{R})$. However, as there are no eigenvalues of $A(\lambda)$ inside $D(0,\tilde{R})$, $T(\lambda)=A(\lambda)+B(\lambda)$ does not have any eigenvalues inside $D(0,\tilde{R})$. Thus, for any eigenvalue $\lambda_{0}$ of $T(\lambda)$ we have $\tilde{R}\leq|\lambda_{0}|$, thereby giving a lower bound as required. ∎
2.4. Bound on the eigenvalues of $T(\lambda)$ using polynomials
Let us now consider yet another well known technique that is used to find bounds on the eigenvalues of matrix polynomials. The idea is to convert the rational matrix $T(\lambda)$ given in Equation (2.2) into a matrix polynomial by multiplying by $\prod_{i=1}^{m}(\lambda-\alpha_{i})$. That is,
$\prod_{i=1}^{m}(\lambda-\alpha_{i})T(\lambda)=P(\lambda),$ (2.5)
a matrix polynomial. One can verify that the set of eigenvalues of $T(\lambda)$ is contained in the set of eigenvalues of $P(\lambda)$. While there are many techniques in the literature to determine the eigenvalue location of matrix polynomials, we restrict ourselves to only one such method due to Higham and Tisseur (Lemma $3.1$, [11]), stated below.
Lemma 2.8.
Let $P(\lambda)=\displaystyle\sum_{i=0}^{m}A_{i}\lambda^{i}$ be an $n\times n$ matrix polynomial with $A_{m}$, $A_{0}$ being nonsingular. Define two scalar polynomials associated with $P(\lambda)$ as follows:
$u(\lambda)=||A_{m}^{-1}||^{-1}\lambda^{m}-||A_{m-1}||\lambda^{m-1}-\cdots-||A_{0}||$, $l(\lambda)=||A_{m}||\lambda^{m}+\cdots+||A_{1}||\lambda-||A_{0}^{-1}||^{-1}$,
where |||||$\cdot$||| | ⋅ | | is any induced matrix norm. Then every eigenvalue $\lambda_{0}$ of $P$ satisfies
$r\leq|\lambda_{0}|\leq R$,
where $R$ and $r$ are the unique real positive root of $u(\lambda)$ and $l(\lambda)$ respectively.
When $m$ is large, it is difficult to determine the coefficients of the matrix polynomial $P(\lambda)$ described in the previous paragraph. We, therefore, restrict ourselves to the case when $m=1$. This is mainly for the sake of comparison and the exact proof carries over to arbitrary $m$.
Theorem 2.9.
Let $T(\lambda)=-B_{0}+I\lambda+\frac{B_{1}}{\lambda-\alpha}$, where $\alpha$ is a complex number. Then for any eigenvalue $\lambda_{0}$ of $T(\lambda)$, $|\lambda_{0}|\leq R$, where $R$ is the unique positive root of the polynomial $u(\lambda)=\lambda^{2}-||B_{0}+\alpha I||\lambda-||\alpha B_{0}+B_{1}||$.
Proof.
Let $P(\lambda)=(\lambda-\alpha)T(\lambda)=I\lambda^{2}-(B_{0}+\alpha I)\lambda+(\alpha B_{0}+B_{1})$. Note that if $\lambda_{0}\in\mathbb{C}$ is an eigenvalue of $T(\lambda)$ then $\lambda_{0}$ is an eigenvalue of $P(\lambda)$ also. We deduce the desired conclusion from Lemma 2.8. ∎
2.5. Estimation of bound on the largest root of scalar rational function $q(x)$
Let $T(\lambda)$ be as in Equation (2.2). By Theorem 2.5 we know that the largest real root $R$, of the scalar rational function $q(x)$ given in Equation (2.3) is an upper bound on the moduli of the eigenvalues of $T(\lambda)$. $q(x)$ being a $1\times 1$ rational matrix, its roots are contained in the set of eigenvalues of the matrix $C_{q}=\begin{bmatrix}|\alpha_{1}|&0&\cdots&0&-1\\ 0&|\alpha_{2}|&\cdots&0&-1\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&|\alpha_{m}|&-1\\ -||B_{1}||&-||B_{2}||&\cdots&-||B_{m}||&||B_{0}||\end{bmatrix}$ of size $(m+1)\times(m+1)$. We now give a bound on $R$ using the numerical radius of $C_{q}$. We begin with a few notations. Given $A\in M_{n}(\mathbb{C})$ the numerical range and the numerical radius of $A$ are denoted by $W(A)$ and $w(A)$ respectively and are defined as $W(A):=\{x^{*}Ax:x\in\mathbb{C}^{n}\,\text{and}\,||x||=1\}$ and $w(A)=\sup\{|\lambda|:\lambda\in W(A)\}$. If $\rho(A)$ denotes the spectral radius of $A$, then for any eigenvalue $\mu_{0}$ of $A$ we have $|\mu_{0}|\leq\rho(A)\leq w(A)$.
We use the following lemma to estimate a bound on $R$.
Lemma 2.10 ([1], Lemma $3$).
Let $A\in M_{k}(\mathbb{C}),B\in M_{k,s}(\mathbb{C}),C\in M_{s,k}(\mathbb{C})$ and $D\in M_{s}(\mathbb{C})$, and let $K=\begin{bmatrix}A&B\\ C&D\end{bmatrix}$. Then $w(K)\leq\displaystyle\frac{1}{2}\Big{(}w(A)+w(D)+\sqrt{(w(A)-w(D))^{2}+4w^{2}(K_{0})}\Big{)}$, where $K_{0}=\begin{bmatrix}0&B\\ C&0\end{bmatrix}$.
Theorem 2.11.
Let $q(x)$ be a rational function as in Equation (2.3), and $R$ be the largest root of $q(x)$. Then
$R\leq\displaystyle\frac{1}{2}\Bigg{(}|\alpha_{m}|+||B_{0}||+\sqrt{(|\alpha_{m}|-||B_{0}||)^{2}+4\displaystyle\sum_{i=1}^{m}||B_{i}||}\Bigg{)}$.
Proof.
Let us write $C_{q}=\begin{bmatrix}\begin{array}[]{c c c c|c}|\alpha_{1}|&0&\cdots&0&-1\\ 0&|\alpha_{2}|&\cdots&0&-1\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&|\alpha_{m}|&-1\\ \hline\cr-||B_{1}||&-||B_{2}||&\cdots&-||B_{m}||&||B_{0}||\end{array}\end{bmatrix}=\begin{bmatrix}A&B\\ C&D\end{bmatrix}$. Then $w(A)=|\alpha_{m}|$ and $w(D)=||B_{0}||$. Let $K_{0}:=\begin{bmatrix}0&\cdots&0&-1\\ 0&\cdots&0&-1\\ \vdots&\ddots&\vdots&\vdots\\ 0&\cdots&0&-1\\ -||B_{1}||&\cdots&-||B_{m}||&0\end{bmatrix}$. Since $K_{0}$ is a real symmetric matrix, $\rho(K_{0})=w(K_{0})$. To determine $\rho(K_{0})$ we first prove that the characteristic polynomial of $K_{0}$ is $t(\lambda)=(-1)^{m+1}\lambda^{m+1}+(-1)^{m}\displaystyle\sum_{i=1}^{m}||B_{i}||\lambda^{m-1}$. We prove this by induction on $m$, the number of poles. If $m=1$, then $K_{0}=\begin{bmatrix}0&-1\\ -||B_{1}||&0\end{bmatrix}$. Therefore, $t(\lambda)=\begin{vmatrix}-\lambda&-1\\ -||B_{1}||&-\lambda\end{vmatrix}=\lambda^{2}-||B_{1}||$. This proves the claim for $m=1$. Assume that the claim is true when there are $m$ poles; that is, the characteristic polynomial of $K_{0}=\begin{bmatrix}0&\cdots&0&-1\\ 0&\cdots&0&-1\\ \vdots&\ddots&\vdots&\vdots\\ 0&\cdots&0&-1\\ -||B_{1}||&\cdots&-||B_{m}||&0\end{bmatrix}$ is $t(\lambda)=(-1)^{m+1}\lambda^{m+1}+(-1)^{m}\sum_{i=1}^{m}||B_{i}||\lambda^{m-1}$. Now consider the case when there are $m+1$ poles. In this case $K_{0}=\begin{bmatrix}0&\cdots&0&-1\\ 0&\cdots&0&-1\\ \vdots&\ddots&\vdots&\vdots\\ 0&\cdots&0&-1\\ -||B_{1}||&\cdots&-||B_{m+1}||&0\end{bmatrix}$. The characteristic polynomial of $K_{0}$ is $t(\lambda)=\begin{vmatrix}-\lambda&0&\cdots&0&-1\\ 0&-\lambda&\cdots&0&-1\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&-\lambda&-1\\ -||B_{1}||&-||B_{2}||&\cdots&-||B_{m+1}||&-\lambda\end{vmatrix}=(-\lambda)\begin{vmatrix}-\lambda&\cdots&0&-1\\ \vdots&\ddots&\vdots&\vdots\\ 0&\cdots&-\lambda&-1\\ -||B_{2}||&\cdots&-||B_{m+1}||&-\lambda\end{vmatrix}+(-1)^{m+2}\begin{vmatrix}0&-\lambda&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&-\lambda\\ -||B_{1}||&-||B_{2}||&\cdots&-||B_{m+1}||\end{vmatrix}$. By induction we have,
$\displaystyle t(\lambda)$ $=(-\lambda)\Big{(}(-1)^{m+1}\lambda^{m+1}+(-1)^{m}\sum_{i=2}^{m+1}||B_{i}||\lambda^{m-1}\Big{)}+(-1)^{m+1}||B_{1}||\lambda^{m}$
$=(-1)^{m+2}\lambda^{m+2}+(-1)^{m+1}\sum_{i=2}^{m+1}||B_{i}||\lambda^{m}+(-1)^{m+1}||B_{1}||\lambda^{m}$
$=(-1)^{m+2}\lambda^{m+2}+(-1)^{m+1}\sum_{i=1}^{m+1}||B_{i}||\lambda^{m}.$
Hence, by mathematical induction if $q(x)$ has $m$ number of poles then the characteristic polynomial of $K_{0}$ is $t(\lambda)=(-1)^{m+1}\lambda^{m+1}+(-1)^{m}\sum_{i=1}^{m}||B_{i}||\lambda^{m-1}$. This implies $\rho(K_{0})=w(K_{0})=\displaystyle\sqrt{\sum_{i=1}^{m}||B_{i}||}$. Therefore, by Lemma 2.10 we have $R\leq\displaystyle\frac{1}{2}\Bigg{(}|\alpha_{m}|+||B_{0}||+\sqrt{(|\alpha_{m}|-||B_{0}||)^{2}+4\displaystyle\sum_{i=1}^{m}||B_{i}||}\Bigg{)}$. ∎
Remark 2.12.
If $\lambda_{0}$ is an eigenvalue of $T(\lambda)$ as given in Equation (2.2), then by above theorem
$|\lambda_{0}|\leq\displaystyle\frac{1}{2}\Bigg{(}|\alpha_{m}|+||B_{0}||+\sqrt{(|\alpha_{m}|-||B_{0}||)^{2}+4\displaystyle\sum_{i=1}^{m}||B_{i}||}\Bigg{)}.$
3. Comparison of bounds
In this section, we compare the bounds obtained in Section 2. In general, we cannot determine which method gives a better bound (see Remark 3.3) for arbitrary matrix coefficients. But when the coefficients of rational matrices as given in Equation (2.2) are unitary matrices, we have the following comparison of the bounds given in Theorems 2.3 and 2.5 and Theorems 2.3 and 2.11.
Theorem 3.1.
Let $T(\lambda)=-B_{0}+I\lambda+\frac{B_{1}}{\lambda-\alpha_{1}}+\dots+\frac{B_{m}}{\lambda-\alpha_{m}}$, where $B_{i}$’s are $n\times n$ unitary matrices and $\alpha_{i}$’s are distinct complex numbers. Then the bound given in Theorem 2.5 is better than the bound given in Theorem 2.3.
Proof.
Let $R_{1}$ be the bound for the eigenvalues of $T(\lambda)$ given in Theorem 2.3; that is, $R_{1}=a+|\alpha_{m}|$, where $a=\displaystyle\Bigg{[}\frac{(2m+1)+(4m+1)^{1/2}}{2}\Bigg{]}^{1/2}$. Let $R_{2}$ be the eigenvalue bound given in Theorem 2.5, which is the unique root of the real rational function $q(x)=x-1-\frac{1}{x-|\alpha_{1}|}-\dots-\frac{1}{x-|\alpha_{m}|}$ such that $|\alpha_{i}|<R_{2}$ for all $i=1,2,\dots,m$. If $m=1$ and $\alpha_{1}=0$, then it is easy to verify that $R_{1}=R_{2}$. Therefore, we assume that $m\geq 1$ and $\alpha_{m}\neq 0$. Note that $R_{1}>|\alpha_{m}|$ and $\lim_{x\rightarrow|\alpha_{m}|^{+}}q(x)=-\infty$. Consider
$q(R_{1})$ $=q(a+|\alpha_{m}|)$
$=a+|\alpha_{m}|-1-\frac{1}{a+|\alpha_{m}|-|\alpha_{1}|}-\dots-\frac{1}{a+|\alpha_{m}|-|\alpha_{m}|}$
$=|\alpha_{m}|+\frac{a^{2}-a-1}{a}-\frac{1}{a+|\alpha_{m}|-|\alpha_{1}|}-\dots-\frac{1}{a+|\alpha_{m}|-|\alpha_{m-1}|}$
$>|\alpha_{m}|+\frac{a^{2}-a-1}{a}-\frac{1}{a}-\dots-\frac{1}{a}$
$=|\alpha_{m}|+\frac{a^{2}-a-1}{a}-\frac{m-1}{a}=|\alpha_{m}|+\frac{a^{2}-a-m}{a}.$
It is easy to verify that $a^{2}-a-m=0$ and therefore, $q(R_{1})>|\alpha_{m}|>0$. The intermediate value theorem ensures that $q(x)$ has a root in $(|\alpha_{m}|,R_{1})$. But $R_{2}$ is the only root of $q(x)$ such that $|\alpha_{m}|<R_{2}$. Therefore, $R_{2}\in(|\alpha_{m}|,R_{1})$. Hence, $R_{2}<R_{1}$. ∎
Theorem 3.2.
Let $T(\lambda)=-B_{0}+I\lambda+\frac{B_{1}}{\lambda-\alpha_{1}}+\dots+\frac{B_{m}}{\lambda-\alpha_{m}}$, where $B_{i}$’s are $n\times n$ unitary matrices and $\alpha_{i}$’s are distinct complex numbers. Then the bound given in Theorem 2.11 is better than the bound given in Theorem 2.3.
Proof.
Let $R_{1}=a+|\alpha_{m}|$ be the bound for the eigenvalues of $T(\lambda)$ given in Theorem 2.3, where $a=\displaystyle\Bigg{[}\frac{(2m+1)+(4m+1)^{1/2}}{2}\Bigg{]}^{1/2}$. Let $R_{4}$ be the bound on the eigenvalues of $T(\lambda)$ obtained in Theorem 2.11. Since $B_{i}$’s are unitary matrices, $R_{4}=\displaystyle\frac{1}{2}\Big{(}1+|\alpha_{m}|+\sqrt{(|\alpha_{m}|-1)^{2}+4m}\Big{)}$. If $m=1$ and $\alpha_{1}=0$, then it is easy to check that $R_{1}=R_{4}$. Consider $m\geq 1$ and $\alpha_{m}\neq 0$. Define a scalar rational function $w(x):=x-1-\displaystyle\frac{m}{x-|\alpha_{m}|}$. The only zeros of $w(x)$ are $R_{4}^{\prime}=\displaystyle\frac{1}{2}\Big{(}1+|\alpha_{m}|-\sqrt{(|\alpha_{m}|-1)^{2}+4m}\Big{)}$ and $R_{4}=\displaystyle\frac{1}{2}\Big{(}1+|\alpha_{m}|+\sqrt{(|\alpha_{m}|-1)^{2}+4m}\Big{)}$. By Lemma 2.4, $R_{4}^{\prime}\in(-\infty,|\alpha_{m}|)$ and $R_{4}\in(|\alpha_{m}|,\infty)$. The remaining part of the proof follows as in the previous theorem. ∎
Few remarks are in order. In what follows, we work with the spectral norm.
Remark 3.3.
  1. (1)
    In Theorems 3.1 and 3.2, if the matrices are not unitary, we cannot determine which theorem gives the better bound. For example, consider $T(\lambda)=\displaystyle-B_{0}+I\lambda+\frac{B_{1}}{\lambda-\alpha_{1}}$, where $B_{0}=\begin{bmatrix}0&0\\ 1&0\end{bmatrix}$ and $B_{1}=\begin{bmatrix}1&0\\ 0&0\end{bmatrix}$.
    • a.
      If $\alpha_{1}=0.1$, then the bound given in Theorem 2.2 is $R_{1}=1.51$ and the bound given in Theorem 2.5 is $R_{2}=1.65$. Therefore, $R_{1}<R_{2}$.
    • b.
      If $\alpha_{1}=1$, then the bound given in Theorem 2.2 is $R_{1}=2.41$ and the bound given in Theorem 2.5 is $R_{2}=2$. Therefore, $R_{2}<R_{1}$.
    • c.
      Again, if $\alpha_{1}=0.1$, the bound given in Theorem 2.11 is $R_{4}=1.65$. Therefore, $R_{1}<R_{4}$. If $\alpha=1$ the bound given in Theorem 2.11 is $R_{4}=2$. Hence, $R_{4}<R_{1}$.
  2. (2)
    We cannot say which theorem gives a better bound between Theorem 2.3 and Theorem 2.9, even when the coefficients of $T(\lambda)$ are unitary matrices. For example consider, $T(\lambda)=\displaystyle-B_{0}+I\lambda+\frac{B_{1}}{\lambda-\alpha_{1}}$, where $B_{0}=B_{1}=I_{2}$.
    • a.
      If $\alpha_{1}=1$, then the bound given in Theorem 2.3 is $R_{1}=2.62$ and the bound given in Theorem 2.9 is $R_{3}=1+\sqrt{3}=2.73$. Therefore, we have $R_{1}<R_{3}$.
    • b.
      If and $\alpha_{1}=i$, then the bound given in Theorem 2.3 is $R_{1}=2.62$ and the bound given in Theorem 2.9 is $R_{3}=2.09$. In this case, $R_{3}<R_{1}$.
  3. (3)
    The same phenomenon happens with Theorems 2.5 and 2.9. Consider the same example as in (2).
    • a.
      If $\alpha_{1}=-1.5$, then the bounds given in Theorems 2.5 and 2.9 are $R_{2}=2.28$ and $R_{3}=1$ respectively. Thus, $R_{3}<R_{2}$.
    • b.
      If $\alpha_{1}=1.5$, then the bounds given in Theorems 2.5 and 2.9 are $R_{2}=2.28$ and $R_{3}=3.27$ respectively. In this case, we have $R_{2}<R_{3}$.
  4. (4)
    The bounds obtained in Theorem 2.11 and Theorem 2.9 are also not comparable. Consider the same example given in $(2)$.
    • a.
      If $\alpha_{1}=1$, then bounds obtained in Theorems 2.11 and 2.9 are $R_{4}=2$ and $R_{3}=2.73$ respectively. Therefore, $R_{4}<R_{3}$.
    • b.
      If $\alpha_{1}=-0.5$, then bounds obtained in Theorems 2.11 and 2.9 are $R_{4}=1.78$ and $R_{3}=1$ respectively. Therefore, $R_{3}<R_{4}$.
4. Numerical results
We present two examples of REPs and compare the bounds obtained in the previous sections with bounds that are available in the literature. The first one arises in the finite element discretization of a boundary problem describing the eigenvibration of a string with a load of mass attached by an elastic spring. We refer readers to [4] for details about this particular REP. The second one is chosen randomly.
Example 4.1.
Let
$T(\lambda)v=\Big{(}A-B\lambda+C\frac{\lambda}{\lambda-\alpha}\Big{)}v,$
where $A=\frac{1}{h}\begin{bmatrix}2&-1&&\\ -1&\ddots&\ddots&\\ &\ddots&2&-1\\ &&-1&1\end{bmatrix}$, $B=\frac{h}{6}\begin{bmatrix}4&1&&\\ 1&\ddots&\ddots&\\ &\ddots&4&1\\ &&1&2\end{bmatrix}$ and $C=e_{n}e_{n}^{T}$, are $n\times n$ matrices with $h=\frac{1}{n}$. Note that $B$ is invertible as it is strictly diagonally dominant. Therefore, the above REP is the same as
$\Big{(}-(A+C)B^{-1}+I\lambda-CB^{-1}\frac{\alpha}{\lambda-\alpha}\Big{)}v=0.$
Table 1:
$n$ $\mu$ $R_{1}$ $R_{2}$ $R_{3}$ $R_{4}$
3 94.60 98.46 97.38 99.18 97.38
5 285.16 292.89 291.85 293.74 291.85
10 1184.46 1203.05 1202.03 1203.98 1202.03
100 119984.22 120663.52 120662.49 120664.48 120662.49
1000 11999984.22 12062993.53 12062992.50 12062994.49 12062992.50
10000 1199999984.22 1206267189.08 1206267188.06 1206267190.05 1206267188.06
Table 1. Comparison of bounds on the eigenvalues of $T(\lambda)$ as $n$ varies.
In Table 1, given above, we obtain bounds on eigenvalues of $T(\lambda)$ using the methods given in Section 2. We assume $\alpha=1$ and use the spectral norm for the calculations. Let $n$ denote the size of the rational matrix $T(\lambda)$ and $\mu$ be the maximum of the moduli of the eigenvalues of $T(\lambda)$. Let $R_{1},R_{2},R_{3}$ and $R_{4}$ denote the bounds obtained in Theorems 2.2, 2.5, 2.9 and 2.11 respectively. Table 2 also suggests that for Example 4.1 the bounds obtained using Theorem 2.5 is better than the ones obtained using Theorems 2.2 and 2.9.
In this example, it is interesting to note that the differences among the bounds $R_{1}$, $R_{2}$, $R_{3}$ and $R_{4}$ are of the same order for all $n$ although the spectral radius $\mu$ of $T(\lambda)$ becomes very large for large $n$. Each term in the $R_{i}$’s involves the norm of combination of the product and sum of matrices $A$, $C$ and $B^{-1}$ and these norms depend on the size of the matrices. Therefore, we need to express each term in terms of $n$, the size of the matrices. Though there are formulas to compute the inverse of the symmetric tridiagonal matrix $B$ (see for instance, [10]), it is difficult to determine the norm of matrices involving $B^{-1}$. However, from MATLAB calculations we observe that if $A$, $B$ and $C$ are of size $n\times n$ then we have following approximations :
$\begin{split}||(A+C)B^{-1}+I||&\approx 1+||(A+C)B^{-1}||\approx 12n^{2}:=\alpha,\\ ||CB^{-1}||&\approx 3.6n,\\ ||AB^{-1}||-\beta&\approx 3n,\text{where}\ \beta:=||(A+C)B^{-1}||-||CB^{-1}||,\\ ||E||&\approx 12n^{2}-1.\end{split}$
On substituting these values in $R_{i}$’s we have
$\begin{split}2(R_{3}-R_{4})&=\sqrt{{\alpha}^{2}+4(3n+\beta)}-\sqrt{{\alpha}^{2}-4\beta}\\ &=\sqrt{144n^{4}+12n+48n^{2}-14.4n-4}-\sqrt{144n^{4}-48n^{2}+14.4n+4}\end{split}.$
Multiplying and dividing the above equation by
$\sqrt{144n^{4}+12n+48n^{2}-14.4n-4}+\sqrt{144n^{4}-48n^{2}+14.4n+4}$ we obtain
$\begin{split}2(R_{3}-R_{4})&=\frac{96n^{2}-16.8n-8}{\sqrt{144n^{4}+12n+48n^{2}-14.4n-4}+\sqrt{144n^{4}-48n^{2}+14.4n+4}}\\ \\ &=\frac{n^{2}(96-\frac{16.8}{n}-\frac{8}{n^{2}})}{n^{2}\sqrt{144+\frac{12}{n^{3}}+\frac{48}{n^{2}}-\frac{14.4}{n^{3}}-\frac{4}{n^{4}}}+n^{2}\sqrt{144-\frac{48}{n^{2}}+\frac{14.4}{n^{3}}+\frac{4}{n^{4}}}}\end{split}.$
Now as $n\rightarrow\infty$, $2(R_{3}-R_{4})\rightarrow 4$. Thus, $R_{3}-R_{4}\approx 2$. Using similar techniques it is easy to show $R_{1}-R_{4}\approx 1$ and $R_{3}-R_{1}\approx 1$. In $R_{4}$, rewriting the term inside the square root, we see that $R_{2}=R_{4}$. But in general by Theorem $2.11$, $R_{2}\leq R_{4}$.
To the best of our knowledge, there are no bounds on the moduli of eigenvalues of rational matrices in the literature so far. There are methods to determine approximate eigenvalues in specific regions using iterative methods and rational approximation methods (see for instance, [16], [17]), which are entirely different problems from ours. However, one can convert the rational matrix to a matrix polynomial and use existing results in the literature on matrix polynomials to compare these bounds. We do this and compare our bounds with a bound due to Theorem $2.1$ of [5] and bounds given in [13]. Let $P(\lambda)=(\lambda-1)T(\lambda)=I\lambda^{2}-[(A+C)B^{-1}+I]\lambda+AB^{-1}$. Note that the eigenvalues of $T(\lambda)$ are eigenvalues of $P(\lambda)$. Using item $(4)$ of Theorem $2.1$ in [5], the bound on moduli of eigenvalues of $P(\lambda)$ is the unique positive real root, say $R_{5}$, of the polynomial $p(x)=x^{2}-||(A+C)B^{-1}+I||x-||AB^{-1}||$. But this polynomial is same as the polynomial $u(\lambda)$ in Theorem $2.9$ of our manuscript, hence, $R_{5}=R_{3}$. In the table below, we provide bounds obtained using several theorems and corollaries from [13]. Let us point out that Roy and Bora [15] also study the eigenvalue location of quadratic matrix polynomials and compare their bounds with that of [5]; however, the bounds in [5] are better than that of [15].
Table 2:
Order Theorems/Corollaries of [13]
$n$ 3.2 3.3, 3.6 3.4 3.4.2 3.4.4 ($R_{6}$) 3.4.6 ($R_{7}$)
3 99.24 99.25 99.19 101.13 97.33 98.33
5 293.76 293.77 293.75 295.71 290.37 291.37
10 1203.98 1203.99 1203.98 1205.96 1196.16 1197.17
100 120664.48 120664.48 120664.48 120666.47 120282.67 120283.67
1000 12062994.49 12062994.49 12062994.49 12062996.48 12025685.14 12025686.14
10000 1206267190.05 1206267190.05 1206267190.05 1206267192.04 1202544209.92 1202544210.92
Table 2. Bounds obtained using Theorems/Corollaries of [13]
Table 2 suggests that for large $n$, bounds obtained using Corollaries $3.4.4$ and $3.4.6$, say $R_{6}$ and $R_{7}$, are better than all our bounds. For large $n$ we have the following comparison: $R_{6}<R_{7}<R_{2}=R_{4}<R_{1}<R_{3}=R_{5}$. However, all other bounds obtained from [13] are greater than or equal to $R_{3}$. But as mentioned in the beginning of this section, one cannot in general determine which method will give the best possible bound. For instance, consider the following example, where $R_{6}$, $R_{7}$ are not better bounds when compared to ours.
Example 4.2.
Let $T(\lambda)=-B_{0}+I\lambda+\displaystyle\frac{B_{1}}{\lambda-1}$, where $I$ is an $n\times n$ identity matrix and $B_{0}=\text{diag}(1,\ldots,1,2)$, $B_{1}=\text{diag}(0,\ldots,0,-1)$ are $n\times n$ diagonal matrices. Therefore, $T(\lambda)=\text{diag}(\lambda-1,\ldots,\lambda-1,\lambda-2-\frac{1}{\lambda-1})$. It is easy to check that eigenvalues of $T(\lambda)$ are $\frac{3-\sqrt{5}}{2}$ and $\frac{3+\sqrt{5}}{2}\approx 2.62$, with the largest being $2.62$. Since $1$ is a pole of $T(\lambda)$, it is not an eigenvalue. Since entries of $B_{0}$ and $B_{1}$ do not involve $n$, the bounds obtained due to [5], [13] and the ones we have derived are constant for any $n$. The following table gives these bounds for any $n$.
Table 3:
Our bounds [5] Theorems/Corollaries of [13]
$R_{1}$ $R_{2}$ $R_{3}$ $R_{4}$ $R_{5}$ 3.2 3.3, 3.6 3.4 3.4.2 3.4.4 ($R_{6}$) 3.4.6 ($R_{7}$)
3.41 2.62 3.30 2.62 3.30 3.79 4.00 3.41 5.00 3.37 4.32
Table 3. Comparison of bounds
Note that $R_{2}=R_{4}<R_{3}=R_{5}<R_{6}<R_{1}$, while the other bounds obtained from [13] are greater than or equal to $R_{1}$. Therefore, in this example, bounds $R_{2}$, $R_{4}$, $R_{3}$ are better bounds. We end by pointing out that calculating the coefficients of the matrix polynomial that is obtained from the rational matrix is more challenging and not easy to compute, especially when there are more number of poles. This will in turn make it difficult to obtain bounds using this method.
5. Concluding Remarks
We have derived bounds on the eigenvalues of rational matrices using three different methods and have compared the same for certain class of matrix coefficients. Numerical calculations suggest that, in general, the methods discussed in Section 2 are not comparable. However, if the coefficients are unitary, bounds obtained using Theorems 2.5 and 2.11 are better than the one given by Theorem 2.3. We have also considered numerical examples of REPs and compared our bounds with those that exist in the literature, by converting the rational matrix to a matrix polynomial. There are examples where numerical computations suggest that the bounds obtained using Theorem 2.5 might be better than the ones obtained from the other methods discussed. One can consider a monic matrix polynomial of higher degree in place of the pencil $-B_{0}+I\lambda$ in Equation (2.1) and extend the results given in this manuscript. Some work in this direction has been done and the details are available in arXiv:2306.14776.
Data availability: No data was used for the research done in this article.
Declarations: The authors declare that there is no conflict of interest in this work.
Acknowledgements: Pallavi Basavaraju and Shrinath Hadimani are grateful for the financial assistance provided by the Council of Scientific and Industrial Research (CSIR) and the University Grants Commission (UGC), both departments of the Indian government.