### v6

# Spectral bounds for certain special type of rational matrices

Abstract.
The aim of this manuscript is to derive bounds on the moduli of eigenvalues of special type of rational matrices of the form $T(\lambda)=\displaystyle-B_{0}+I\lambda+\frac{B_{1}}{\lambda-\alpha_{1}}+\dots+\frac{B_{m}}{\lambda-\alpha_{m}}$, where $B_{i}$’s are $n\times n$ complex matrices and $\alpha_{i}$’s are distinct complex numbers, using the following methods: $(1)$ an upper bound is obtained using the Bauer-Fike theorem for complex matrices on an associated block matrix $C_{T}$ of the given rational matrix $T(\lambda)$, $(2)$ a lower bound is obtained in terms of a zero of a scalar real rational function $p(x)$ associated with $T(\lambda)$, using Rouché’s theorem for matrix-valued functions and $(3)$ an upper bound is also obtained using a numerical radius inequality for a block matrix $C_{q}$ associated with another scalar real rational function $q(x)$ corresponding to $T(\lambda)$. These bounds are compared when the coefficients are unitary matrices. Numerical examples are given to illustrate the results obtained.
Key words and phrases:
Rational matrices; rational eigenvalue problems; spectral bounds for rational matrices; matrix-valued functions; matrix polynomials; numerical radius.
2010 Mathematics Subject Classification:
15A18, 15A42, 15A22, 93B18, 47A56
1. Introduction
An $n\times n$ rational matrix denoted by $T(\lambda)$, is one whose entries are complex rational functions. The rational eigenvalue problem is to determine a complex number $\lambda_{0}$ and an $n\times 1$ nonzero vector $v$ such that $T(\lambda_{0})v=0$. The scalar $\lambda_{0}$ is called an eigenvalue of $T(\lambda)$. Rational eigenvalue problems, abbreviated henceforth as REPs, are an important class of nonlinear eigenvalue problems that arise in applications to science and engineering. REPs arise for instance, in applications to computing damped vibration modes of an acoustic fluid confined in a cavity [8], describing the eigenvibration of a string with a load of mass attached by an elastic string [5] and in application to photonic crystals [13], to name a few. Readers may refer to [4], [12], [17], [19] and [20] and the references cited therein for some recent work on REPs. Determining the exact eigenvalues of rational matrices presents a significant challenge, and hence they are approximated using iterative methods [17]. Consequently, establishing bounds on eigenvalues is crucial for making an initial guess, which influences the convergence rate of the iteration. A rational matrix, where each entry is a scalar polynomial, is referred to as a matrix polynomial. There are several spectral bounds for matrix polynomials depending on norms of coefficients and roots of associated scalar polynomials (see for instance, [7] and [14]). For rational matrices, one approach to estimate/compute spectral bounds is to convert the rational matrix to a matrix polynomial and use existing results. In practice, it is difficult to determine the coefficients of the matrix polynomial that comes out from this technique. Among the ones available in the literature, there is no easy or the best way to determine the location of eigenvalues of rational matrices using matrix polynomials. The purpose of this work is to derive bounds on the moduli of eigenvalues of certain special type of rational matrices. We provide bounds that can be calculated with a small computational effort.
We work either over the field $\mathbb{C}$ of complex numbers or over the field $\mathbb{R}$ of real numbers. The vector space of $n\times n$ matrices over $\mathbb{C}$ (respectively, $\mathbb{R}$) is denoted by $M_{n}(\mathbb{C})$ (respectively, $M_{n}(\mathbb{R})$). $||\cdot||_{2}$ denotes the spectral norm of a square matrix. The condition number of a square matrix $A$ is denoted by $\kappa(A)$ is defined as $\kappa(A)=\begin{cases}||A||||A^{-1}||&\text{if $A$ is invertible}\\ \infty&\text{otherwise}\end{cases}$, where |||||$\cdot$||| | ⋅ | | is any matrix norm.
The organization of the manuscript is as follows. Section 2.1 contains a brief introduction to rational matrices. This is followed by deriving a bound on the moduli of eigenvalues of rational matrices of special type using an associated block matrix (see Section 2.1 for the definition and Section 2.2 for details). We then derive a lower bound on the moduli of eigenvalues using roots of a real rational function (see Section 2.3 for details). Section 2.4 concerns deriving a bound using scalar polynomials and the same is done using numerical radius in Section 2.5. These bounds are compared in Section 3. Numerical illustrations are given in Section 4. The computations are done using Matlab.
2. Main results
The main results are presented in this section. We start with preliminaries on rational matrices. In the subsections that follow, we derive various bounds on the moduli of eigenvalues of rational matrices.
2.1. Rational matrices
Any $n\times n$ rational matrix can be expressed in the form $T(\lambda)=P(\lambda)-\sum_{i=1}^{k}\frac{s_{i}(\lambda)}{q_{i}(\lambda)}E_{i}$, where $P(\lambda)=\displaystyle\sum_{j=0}^{d}A_{j}\lambda^{j}$ is an $n\times n$ matrix polynomial of degree $d$, $s_{i}(\lambda)$ and $q_{i}(\lambda)$ are scalar polynomials of degree $n_{i}$ and $d_{i}$ respectively, and $A_{j}$’s, $E_{i}$’s $\in M_{n}(\mathbb{C})$. Rational matrices are often known as rational matrix functions or matrix rational functions in the literature. An $n\times n$ rational matrix $T(\lambda)$ is said to be regular if its determinant does not vanish identically. The rational eigenvalue problem (REP) is to find a scalar $\lambda_{0}\in\mathbb{C}$ and a nonzero vector $v\in\mathbb{C}^{n}$ such that $T(\lambda_{0})v=0$, with $T(\lambda)$ being regular and $T(\lambda_{0})$ bounded (that is, $T(\lambda_{0})$ has finite entries). The scalar $\lambda_{0}$ so obtained is called an eigenvalue of $T(\lambda)$ and the vector $v$ is called an eigenvector of $T(\lambda)$ corresponding to the eigenvalue $\lambda_{0}$. Note that if $T(\lambda)=B$, where $B$ is a nonsingular matrix, then no complex number is an eigenvalue for $T(\lambda)$.
The nonlinear eigenvalue problem (abbreviated as NEP) seeks to find a scalar $\lambda_{0}$ and a nonzero vector $v\in\mathbb{C}^{n}$ satisfying $G(\lambda_{0})v=0$, where $G(\lambda)$ is a regular matrix-valued function ($G(\lambda)$ is square and its determinant does not vanish identically), and each entry of $G(\lambda_{0})$ is bounded. The majority of the nonlinear eigenvalue problems in applications are of the form, $G(\lambda)=-B_{0}+A_{0}\lambda+A_{1}f_{1}(\lambda)+\dots+A_{p}f_{p}(\lambda)$, where $f_{i}:\Omega\rightarrow\mathbb{C}$ are analytic functions and $\Omega$ is a region in $\mathbb{C}$. In [19], to study the eigenvalue problem of $G(\lambda)$ authors consider the $surrogate$ problem, $T(\lambda)v=\Big{(}-B_{0}+A_{0}\lambda+\displaystyle\sum_{i=1}^{m}\frac{B_{i}}{\lambda-\alpha_{i}}\Big{)}v=0$, where $\alpha_{i}$’s are distinct complex numbers. This is achieved by approximating $f_{i}$’s by a rational function of the form $r_{i}(\lambda)=\displaystyle\sum_{i=1}^{m}\frac{\sigma_{ij}}{\lambda-\alpha_{i}}$. The $\alpha_{i}$’s are the same for all functions $r_{i}$’s. This motivates us to study the rational matrices of the form
$T(\lambda)=-B_{0}+A_{0}\lambda+\frac{B_{1}}{\lambda-\alpha_{1}}+\dots+\frac{B_{m}}{\lambda-\alpha_{m}},$ (2.1)
where $B_{i}$’s are $n\times n$ matrices and $\alpha_{i}$’s are distinct complex numbers. Since we are interested in finding bounds on the eigenvalues of $T(\lambda)$ we assume $A_{0}$ to be nonsingular. We assume $A_{0}=I$, the identity matrix as one can multiply Equation (2.1) by $A_{0}^{-1}$. We, thus, consider rational matrices of the form
$T(\lambda)=\displaystyle-B_{0}+I\lambda+\frac{B_{1}}{\lambda-\alpha_{1}}+\frac{B_{2}}{\lambda-\alpha_{2}}+\cdots+\frac{B_{m}}{\lambda-\alpha_{m}},$ (2.2)
where $\alpha_{i}$’s are distinct complex numbers ordered $|\alpha_{1}|<|\alpha_{2}|<\cdots<|\alpha_{m}|$ and the $B_{i}$’s are $n\times n$ complex matrices. It turns out that many rational eigenvalue problems can be converted to this form (see for instance, Example 4.3). Note that for $T(\lambda)$ given in Equation (2.2), the REP can be converted to a linear eigenvalue problem, $P(\lambda)v=0$, where $P(\lambda)=I\lambda-C_{T}$ with $C_{T}=\begin{bmatrix}\alpha_{1}I&0&\cdots&0&-I\\ 0&\alpha_{2}I&\cdots&0&-I\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&\alpha_{m}I&-I\\ B_{1}&B_{2}&\cdots&B_{m}&B_{0}&\end{bmatrix}$ of size $(m+1)n\times(m+1)n$. $P(\lambda)$ is also a polynomial system matrix of $T(\lambda)$ with the state matrix $A(\lambda)=\text{diag}\big{(}(\lambda-\alpha_{1})I,\ldots,(\lambda-\alpha_{m})I\big{)}$ (see [3], [12] for details). Interestingly, the linearization of $T(\lambda)$ given in [2] using Fiedler matrices is the same as $P(\lambda)$. Thus, corresponding to $T(\lambda)$ we associate the block matrix $C_{T}$. Note that the eigenvalues of $T(\lambda)$ are also eigenvalues of $C_{T}$. Moreover, if all the $B_{i}$’s are nonsingular, then $T(\lambda)$ and $C_{T}$ have the same eigenvalues (see [12] for details). It is easy to verify that if one of the coefficients $B_{i}$ is singular, then the corresponding pole $\alpha_{i}$ is an eigenvalue of $C_{T}$. Therefore, $T(\lambda)$ has at most $(m+1)n$ eigenvalues.
In [9], the author proves certain perturbation results for eigenvalues of matrix polynomials that are analogous to the Bauer-Fike theorem for complex matrices. The author considers a matrix polynomial $P(\lambda)$ and perturbs the coefficient matrices to get another matrix polynomial $\widetilde{P}(\lambda)$. Further, the author defines the spectral variation between the eigenvalues of $P(\lambda)$ and $\widetilde{P}(\lambda)$ and derive Bauer-Fike type results on the spectral variation using a Jordan triplet of $P(\lambda)$. Similar results were studied in [10] for periodic pairs of matrices. Eigenvalue perturbation theory for homogeneous matrix polynomials can also be found in [11]. Note that one can obtain a rational matrix $\widetilde{T}(\lambda)$ by perturbing the coefficient matrices of the given rational matrix $T(\lambda)$. However, the block companion matrix corresponding to $\widetilde{T}(\lambda)$ and a matrix obtained by perturbing the entries of the block companion matrix $C_{T}$ of $T(\lambda)$ are not the same in general. We do not study perturbation results for rational matrices in this manuscript. Instead, we give a region that contains the eigenvalues of a rational matrix $T(\lambda)$.
2.2. Bounds on the eigenvalues of $T(\lambda)$ using Bauer-Fike theorem
One of the well known results in the perturbation theory of the eigenvalue of a diagonalizable matrix is due to Bauer and Fike (Theorem 6.3.2, [15]). We make use of this result to find a bound for the moduli of eigenvalues of a rational matrix, given in Equation (2.2). We state the Bauer-Fike theorem below.
Theorem 2.1.
Let $A\in M_{n}(\mathbb{{C}})$ be diagonalizable, and suppose that $A=S\Lambda S^{-1}$, in which $S$ is nonsingular and $\Lambda$ is diagonal. Let $E\in M_{n}(\mathbb{C})$ and |||||$\cdot$||| | ⋅ | | be a matrix norm on $M_{n}(\mathbb{C})$ that is induced by an absolute norm on $\mathbb{C}^{n}$. If $\hat{\lambda}$ is an eigenvalue of $A+E$, then there is an eigenvalue $\lambda$ of $A$ such that $|\hat{\lambda}-\lambda|\leq\kappa(S)||E||$ in which $\kappa(\cdot)$ is the condition number with respect to the matrix norm.
The block matrix $C_{T}$ associated with the rational matrix given in Equation (2.2) can be expressed as $C_{T}=A+E$, where $A=\text{diag}(\alpha_{1}I,\ldots,\alpha_{m}I,0)$ and $E=C_{T}-A$. Since $A$ is diagonalizable, the following result is an easy consequence of the Bauer-Fike theorem.
Theorem 2.2.
Let $T(\lambda)$ be as in Equation (2.2), with an eigenvalue $\lambda_{0}$. Then $|\lambda_{0}|\leq||E||+|\alpha_{m}|$, where $E=\begin{bmatrix}0&\cdots&0&-I\\ \vdots&\ddots&\vdots&\vdots\\ 0&\cdots&0&-I\\ B_{1}&\cdots&B_{m}&B_{0}\\ \end{bmatrix}$ and |||||$\cdot$||| | ⋅ | | is any matrix norm induced by an absolute norm on $\mathbb{C}^{n}$.
In particular, if we assume that the $B_{i}$’s are unitary matrices and the induced norm is the spectral norm, then we get a bound which depends only on the number of poles of $T(\lambda)$ and their moduli. We state this below.
Theorem 2.3.
Let $T(\lambda)=-B_{0}+I\lambda+\frac{B_{1}}{\lambda-\alpha_{1}}+\cdots+\frac{B_{m}}{\lambda-\alpha_{m}}$, where $B_{i}$’s are $n\times n$ unitary matrices and $\alpha_{i}$’s are distinct complex numbers. Then $|\lambda_{0}|\leq\displaystyle\Bigg{\{}\frac{(2m+1)+(4m+1)^{1/2}}{2}\Bigg{\}}^{1/2}+|\alpha_{m}|$ for any eigenvalue $\lambda_{0}$ of $T(\lambda)$.
Proof.
By Theorem 2.2, we have $|\lambda_{0}|\leq||E||_{2}+|\alpha_{m}|$. To find $||E||_{2}$, consider $EE^{\ast}=\begin{bmatrix}I&\cdots&I&-B_{0}^{*}\\ \vdots&\ddots&\vdots&\vdots\\ I&\cdots&I&-B_{0}^{*}\\ -B_{0}&\cdots&-B_{0}&(m+1)I\end{bmatrix}$. Consider the matrix $Q(\lambda)$, defined by $Q(\lambda):=\begin{bmatrix}I&0&\cdots&0&0\\ -(I-I\lambda)^{-1}&I&\cdots&0&0\\ -(2I-I\lambda)^{-1}&-(2I-I\lambda)^{-1}&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ -\big{(}(m-1)I-I\lambda\big{)}^{-1}&-\big{(}(m-1)I-I\lambda\big{)}^{-1}&\cdots&I&0\\ B_{0}(mI-I\lambda)^{-1}&B_{0}(mI-I\lambda)^{-1}&\cdots&B_{0}(mI-I\lambda)^{-1}&I\\ \end{bmatrix}$. Observe that $\text{det}Q(\lambda)=1$ and $Q(\lambda)(EE^{*}-I\lambda)$ is a block upper triangular matrix with diagonal blocks $I-I\lambda,I-I\lambda-(I-I\lambda)^{-1},\dots,I-I\lambda-(m-1)\big{(}(m-1)I-I\lambda\big{)}^{-1}$ and $(m+1)I-I\lambda-m(mI-I\lambda)^{-1}$ respectively. Therefore, $\text{det}(EE^{*}-I\lambda)$ equals $\text{det}(Q(\lambda)(EE^{*}-I\lambda))$, which in turn is the product of the determinant of these diagonal blocks. On simplifying, we get
$\text{det}(EE^{*}-I\lambda)$
$=(1-\lambda)^{n}\Big{(}(1-\lambda)-\frac{1}{1-\lambda}\Big{)}^{n}\cdots$
$\hskip 85.35826pt\cdots\Big{(}(1-\lambda)-\frac{(m-1)}{(m-1)-\lambda}\Big{)}^{n}\Big{(}(m+1-\lambda)-\frac{m}{m-\lambda}\Big{)}^{n}$
$=(1-\lambda)^{n}\frac{\lambda^{n}(\lambda-2)^{n}}{(1-\lambda)^{n}}\cdots\frac{\lambda^{n}(\lambda-m)^{n}}{(m-1-\lambda)^{n}}\frac{(\lambda^{2}-(2m+1)\lambda+m^{2})^{n}}{(m-\lambda)^{n}}$
$=(-1)^{(m-1)n}\lambda^{(m-1)n}\big{(}\lambda^{2}-(2m+1)\lambda+m^{2}\big{)}^{n}.$
Thus, the eigenvalues of $EE^{*}$ are 00 and $\frac{(2m+1)\pm(4m+1)^{1/2}}{2}$, from which we see that $||E||_{2}=\displaystyle\Bigg{\{}\frac{(2m+1)+(4m+1)^{1/2}}{2}\Bigg{\}}^{1/2}$. ∎
2.3. Bounds on the eigenvalues of $T(\lambda)$ using rational functions.
Another method for determining bounds on the eigenvalues is to use norms of the coefficient matrices of $T(\lambda)$ and define a rational function whose roots are bounds for the eigenvalues of $T(\lambda)$. In [4], the authors exploit this technique to derive (only) an upper bound on the set of moduli of eigenvalues of a general rational matrix. We state this below (Theorem 2.4) for rational matrices of the form given in Equation 2.2, for the sake of comparison.
Theorem 2.4.
Let $T(\lambda)$ be as given in Equation (2.2), with an eigenvalue $\lambda_{0}$. Define a real rational function associated with $T(\lambda)$ as follows:
$q(x)=x-||B_{0}||-\frac{||B_{1}||}{x-|\alpha_{1}|}-\dots-\frac{||B_{m}||}{x-|\alpha_{m}|},$ (2.3)
where |||||$\cdot$||| | ⋅ | | is any induced matrix norm. Then $|\lambda_{0}|\leq R$, where $R$ is a real positive root of $q(x)$ such that $|\alpha_{i}|<R$ for all $i=1,2,\cdots,m$.
In Theorem 2.7, we derive a lower bound on the set of moduli of eigenvalues of a rational matrix as given in Equation 2.2 that satisfies certain additional assumptions. We do this by associating a real rational function, whose positive root gives a lower bound for the eigenvalues of rational matrices (satisfying additional assumptions). We make use of a Rouché type theorem for analytic matrix-valued functions (Theorem $2.3$ of [7]) and a lemma, whose proof is a simple consequence of the intermediate value theorem. We state them in order of preference.
Theorem 2.5.
Let $A,B:G\rightarrow\mathbb{C}^{n\times n}$, where $G$ is an open connected subset of $\mathbb{C}$, be analytic matrix-valued functions. Assume that $A(\lambda)$ is invertible on the simple closed curve $\gamma\subseteq G$. If $||A(\lambda)^{-1}B(\lambda)||<1$ for all $\lambda\in\gamma$, then $A+B$ and $A$ have the same number of eigenvalues inside $\gamma$, counting multiplicities. The norm is the matrix norm induced by any norm on $\mathbb{C}^{n}$.
Lemma 2.6.
Let $r(x)=x-a_{0}-\frac{a_{1}}{x-b_{1}}-\dots-\frac{a_{m}}{x-b_{m}}$ be a real rational function, where the $a_{i}$’s are positive and $b_{i}$’s are nonnegative real numbers such that $b_{1}<b_{2}<\dots<b_{m}$. Then $r(x)$ has roots $R_{1},R_{2},\dots,R_{m+1}$ such that $R_{1}<b_{1}<R_{2}<b_{2}<\dots<R_{m}<b_{m}<R_{m+1}$.
We now prove the aforesaid theorem that gives a lower bound for the eigenvalues of certain rational matrices.
Theorem 2.7.
Let $T(\lambda)$ be as in Equation (2.2). If $B_{0}$ is invertible and $||B_{0}^{-1}||^{-1}>\displaystyle\frac{||B_{1}||}{|\alpha_{1}|}+\dots+\frac{||B_{m}||}{|\alpha_{m}|}$, then $\tilde{R}\leq|\lambda_{0}|$ for any eigenvalue $\lambda_{0}$ of $T(\lambda)$, where $\tilde{R}$ is the unique positive root of the real rational function $p(x)=x-||B_{0}^{-1}||^{-1}-\frac{||B_{1}||}{x-|\alpha_{1}|}-\dots-\frac{||B_{m}||}{x-|\alpha_{m}|}$ such that $\tilde{R}<|\alpha_{i}|$ for all $i=1,2,\dots,m$.
Proof.
From Lemma 2.6, $p(x)$ has a unique root $\tilde{R}$ such that $\tilde{R}<|\alpha_{1}|$. Note that $p(0)=-||B_{0}^{-1}||^{-1}+\frac{||B_{1}||}{|\alpha_{1}|}+\dots+\frac{||B_{m}||}{|\alpha_{m}|}<0$ by the given hypothesis. Therefore, by the intermediate value theorem $0<\tilde{R}<|\alpha_{1}|$. Let $A(\lambda):=-B_{0}$ and $B(\lambda):=\displaystyle I\lambda+\frac{B_{1}}{\lambda-\alpha_{1}}+\dots+\frac{B_{m}}{\lambda-\alpha_{m}}$. Then $T(\lambda)=A(\lambda)+B(\lambda)$. Taking $G$ to be the disk $D(0,\tilde{R}):=\{z\in\mathbb{C}:|z|<\tilde{R}\}$, we see that $A(\lambda)$ and $B(\lambda)$ are analytic matrix-valued functions on $G$. Since $p(0)<0$ and $p(\tilde{R})=0$, we have $p(x)<0$ for all $0\leq x<\tilde{R}$. Therefore, for all $|\lambda|<\tilde{R}$ we have
$|\lambda|-||B_{0}^{-1}||^{-1}-\frac{||B_{1}||}{|\lambda|-|\alpha_{1}|}-\dots-\frac{||B_{m}||}{|\lambda|-|\alpha_{m}|}<0.$ (2.4)
Now for $|\lambda|<\tilde{R}$, consider
$||B(\lambda)||$ $=\Big{|}\Big{|}I\lambda+\frac{B_{1}}{\lambda-\alpha_{1}}+\dots+\frac{B_{m}}{\lambda-\alpha_{m}}\Big{|}\Big{|}$
$\leq|\lambda|+\frac{||B_{1}||}{|\lambda-\alpha_{1}|}+\dots+\frac{||B_{m}||}{|\lambda-\alpha_{m}|}$
$\leq|\lambda|+\frac{||B_{1}||}{|\alpha_{1}|-|\lambda|}+\dots+\frac{||B_{m}||}{|\alpha_{m}|-|\lambda|}\hskip 28.45274pt(\text{since}\,|\lambda|<|\alpha_{i}|)$
$=|\lambda|-\frac{||B_{1}||}{|\lambda|-|\alpha_{1}|}-\dots-\frac{||B_{m}||}{|\lambda|-|\alpha_{m}|}$
[…]
For any $\epsilon>0$, define $\gamma:=(\tilde{R}-\epsilon)e^{i\theta}$, where $0\leq\theta\leq 2\pi$. Then $||A^{-1}(\lambda)B(\lambda)||<1$ for all $\lambda\in\gamma$. Since $\epsilon>0$ is arbitrary, we see from Theorem 2.5 that the number of eigenvalues of $A(\lambda)$ and $A(\lambda)+B(\lambda)$ are same inside $D(0,\tilde{R})$. However, as there are no eigenvalues of $A(\lambda)$ inside $D(0,\tilde{R})$, $T(\lambda)=A(\lambda)+B(\lambda)$ does not have any eigenvalues inside $D(0,\tilde{R})$. Thus, for any eigenvalue $\lambda_{0}$ of $T(\lambda)$ we have $\tilde{R}\leq|\lambda_{0}|$, thereby giving a lower bound as required. ∎
2.4. Bounds on the eigenvalues of $T(\lambda)$ using polynomials
Let us now consider yet another well known technique that is used to find bounds on the eigenvalues of matrix polynomials. The idea is to convert the rational matrix $T(\lambda)$ given in Equation (2.2) into a matrix polynomial by multiplying by $\prod_{i=1}^{m}(\lambda-\alpha_{i})$. That is,
$\prod_{i=1}^{m}(\lambda-\alpha_{i})T(\lambda)=P(\lambda),$ (2.5)
a matrix polynomial. It is easy to verify that the set of eigenvalues of $T(\lambda)$ is contained in the set of eigenvalues of $P(\lambda)$. While there are many techniques in the literature to determine the eigenvalue location of matrix polynomials, we restrict ourselves to only one such method due to Higham and Tisseur (Lemma $3.1$, [14]).
When $m$ is large, it is difficult to determine the coefficients of the matrix polynomial $P(\lambda)$ described in the previous paragraph. We therefore restrict ourselves to the case when $m=1$. This is mainly for the sake of comparison and the proof carries over verbatim for arbitrary $m$.
Theorem 2.8.
Let $T(\lambda)=-B_{0}+I\lambda+\frac{B_{1}}{\lambda-\alpha}$, where $\alpha$ is a complex number. Then for any eigenvalue $\lambda_{0}$ of $T(\lambda)$, $|\lambda_{0}|\leq R$, where $R$ is the unique positive root of the polynomial $u(\lambda)=\lambda^{2}-||B_{0}+\alpha I||\lambda-||\alpha B_{0}+B_{1}||$.
Proof.
Let $P(\lambda)=(\lambda-\alpha)T(\lambda)=I\lambda^{2}-(B_{0}+\alpha I)\lambda+(\alpha B_{0}+B_{1})$. Note that if $\lambda_{0}\in\mathbb{C}$ is an eigenvalue of $T(\lambda)$, then $\lambda_{0}$ is an eigenvalue of $P(\lambda)$. We deduce the desired conclusion from Lemma $3.1$, [14]. ∎
2.5. Estimation of bounds on the largest root of scalar rational function $q(x)$
Let $T(\lambda)$ be as in Equation (2.2). By Theorem 2.4, we know that the largest real root $R$, of the scalar rational function $q(x)$ given in Equation (2.3) is an upper bound on the moduli of the eigenvalues of $T(\lambda)$. $q(x)$ being a $1\times 1$ rational matrix, its roots are contained in the set of eigenvalues of the matrix $C_{q}=\begin{bmatrix}|\alpha_{1}|&0&\cdots&0&-1\\ 0&|\alpha_{2}|&\cdots&0&-1\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&|\alpha_{m}|&-1\\ -||B_{1}||&-||B_{2}||&\cdots&-||B_{m}||&||B_{0}||\end{bmatrix}$ of size $(m+1)\times(m+1)$. We now give a bound on $R$ using numerical radius inequalities on $C_{q}$. We begin with a few notations. Given $A\in M_{n}(\mathbb{C})$ the numerical range and the numerical radius of $A$ are denoted by $W(A)$ and $w(A)$ respectively and are defined as $W(A):=\{x^{*}Ax:x\in\mathbb{C}^{n}\,\text{and}\,||x||=1\}$ and $w(A)=\sup\{|\lambda|:\lambda\in W(A)\}$. If $\rho(A)$ denotes the spectral radius of $A$, then for any eigenvalue $\mu_{0}$ of $A$ we have $|\mu_{0}|\leq\rho(A)\leq w(A)$.
We use the following lemma to estimate a bound on $R$.
Lemma 2.9 ([1], Lemma $3$).
Let $A\in M_{k}(\mathbb{C}),B\in M_{k,s}(\mathbb{C}),C\in M_{s,k}(\mathbb{C})$ and $D\in M_{s}(\mathbb{C})$, and let $K=\begin{bmatrix}A&B\\ C&D\end{bmatrix}$. Then $w(K)\leq\displaystyle\frac{1}{2}\Big{(}w(A)+w(D)+\sqrt{(w(A)-w(D))^{2}+4w^{2}(K_{0})}\Big{)}$, where $K_{0}=\begin{bmatrix}0&B\\ C&0\end{bmatrix}$.
We first derive a bound on $R$ when the coefficients are unitary matrices.
Theorem 2.10.
Let $q(x)$ be a rational function as in Equation (2.3), and $R$ be the largest root of $q(x)$. If $B_{1},\ldots,B_{m}$ are unitary matrices, then
$R\leq\displaystyle\frac{1}{2}\Bigg{(}|\alpha_{m}|+||B_{0}||_{2}+\sqrt{(|\alpha_{m}|-||B_{0}||_{2})^{2}+4m}\Bigg{)}$.
Proof.
$B_{1},\ldots,B_{m}$ being unitary matrices, $||B_{i}||_{2}=1$ for $1\leq i\leq m$ and $C_{q}=\begin{bmatrix}\begin{array}[]{c c c c|c}|\alpha_{1}|&0&\cdots&0&-1\\ 0&|\alpha_{2}|&\cdots&0&-1\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&|\alpha_{m}|&-1\\ \hline\cr-1&-1&\cdots&-1&||B_{0}||_{2}\end{array}\end{bmatrix}=\begin{bmatrix}A&B\\ C&D\end{bmatrix}$. Then $w(A)=|\alpha_{m}|$ and $w(D)=||B_{0}||_{2}$. Let $K_{0}:=\begin{bmatrix}0&\cdots&0&-1\\ 0&\cdots&0&-1\\ \vdots&\ddots&\vdots&\vdots\\ 0&\cdots&0&-1\\ -1&\cdots&-1&0\end{bmatrix}$. Since $K_{0}$ is a real symmetric matrix, $\rho(K_{0})=w(K_{0})$. It is easy to verify that the characteristic polynomial of $K_{0}$ is $t(\lambda)=(-1)^{m+1}\lambda^{m+1}+(-1)^{m}m\lambda^{m-1}$ so that the eigenvalues of $K_{0}$ are $0,\,\pm\sqrt{m}$. Thus, $\rho(K_{0})=w(K_{0})=\displaystyle\sqrt{m}$. The desired conclusion then follows from Lemma 2.9. ∎
In the above theorem, if $B_{1},B_{2},\ldots,B_{m}$ are arbitrary matrices, then we have the following result.
Theorem 2.11.
Let $q(x)$ be a rational function as in Equation (2.3), and $R$ be the largest root of $q(x)$. Then
$R\leq\displaystyle\frac{1}{2}\Bigg{(}|\alpha_{m}|+||B_{0}||+\sqrt{(|\alpha_{m}|-||B_{0}||)^{2}+4k}\Bigg{)}$,
where $k=\max\left\{m,\displaystyle\sum_{i=1}^{m}||B_{i}||^{2}\right\}$ and |||||$\cdot$||| | ⋅ | | is any induced matrix norm.
Proof.
Consider $C_{q}=\begin{bmatrix}\begin{array}[]{c c c c|c}|\alpha_{1}|&0&\cdots&0&-1\\ 0&|\alpha_{2}|&\cdots&0&-1\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&|\alpha_{m}|&-1\\ \hline\cr-||B_{1}||&-||B_{2}||&\cdots&-||B_{m}||&||B_{0}||\end{array}\end{bmatrix}=\begin{bmatrix}A&B\\ C&D\end{bmatrix}$. Then $w(A)=|\alpha_{m}|$ and $w(D)=||B_{0}||$. Let $K_{0}:=\begin{bmatrix}0&\cdots&0&-1\\ 0&\cdots&0&-1\\ \vdots&\ddots&\vdots&\vdots\\ 0&\cdots&0&-1\\ -||B_{1}||&\cdots&-||B_{m}||&0\end{bmatrix}$. It is easy to verify that $||K_{0}||_{2}=\max\left\{\sqrt{m},\displaystyle\sqrt{\sum_{i=1}^{m}||B_{i}||^{2}}\right\}$. Since $w(K_{0})\leq||K_{0}||_{2}$, the conclusion follows from Lemma 2.9. ∎
Remark 2.12.
If $\lambda_{0}$ is an eigenvalue of $T(\lambda)$ as given in Equation (2.2), then by the above theorem $|\lambda_{0}|\leq\displaystyle\frac{1}{2}\left(|\alpha_{m}|+||B_{0}||+\sqrt{(|\alpha_{m}|-||B_{0}||)^{2}+4\displaystyle k}\right)$, where $k=\max\left\{m,\displaystyle\sum_{i=1}^{m}||B_{i}||^{2}\right\}$.
3. Comparison of bounds
In this section, we compare the bounds obtained in Section 2. In general, we cannot determine which method gives a better bound (see Remark 3.3) for arbitrary matrix coefficients. But when the coefficients of the rational matrices as given in Equation (2.2) are unitary matrices and the norm is the spectral norm, we have the following comparison of the bounds given in Theorems 2.3 and 2.4 and Theorems 2.3 and 2.10.
Theorem 3.1.
Let $T(\lambda)=-B_{0}+I\lambda+\frac{B_{1}}{\lambda-\alpha_{1}}+\dots+\frac{B_{m}}{\lambda-\alpha_{m}}$, where the $B_{i}$’s are $n\times n$ unitary matrices and $\alpha_{i}$’s are distinct complex numbers. Then the bound given in Theorem 2.4 is better than the bound given in Theorem 2.3.
Proof.
Let $R_{1}$ be the bound for the eigenvalues of $T(\lambda)$ given in Theorem 2.3; that is, $R_{1}=a+|\alpha_{m}|$, where $a=\displaystyle\Bigg{[}\frac{(2m+1)+(4m+1)^{1/2}}{2}\Bigg{]}^{1/2}$. Let $R_{2}$ be the eigenvalue bound given in Theorem 2.4, which is the unique root of the real rational function $q(x)=x-1-\frac{1}{x-|\alpha_{1}|}-\dots-\frac{1}{x-|\alpha_{m}|}$ such that $|\alpha_{i}|<R_{2}$ for all $i=1,2,\dots,m$. If $m=1$ and $\alpha_{1}=0$, then it is easy to verify that $R_{1}=R_{2}$. Therefore, we assume that $m\geq 1$ and $\alpha_{m}\neq 0$. Note that $R_{1}>|\alpha_{m}|$ and $\lim_{x\rightarrow|\alpha_{m}|^{+}}q(x)=-\infty$. Consider
$q(R_{1})$ $=q(a+|\alpha_{m}|)$
$=a+|\alpha_{m}|-1-\frac{1}{a+|\alpha_{m}|-|\alpha_{1}|}-\dots-\frac{1}{a+|\alpha_{m}|-|\alpha_{m}|}$
$=|\alpha_{m}|+\frac{a^{2}-a-1}{a}-\frac{1}{a+|\alpha_{m}|-|\alpha_{1}|}-\dots-\frac{1}{a+|\alpha_{m}|-|\alpha_{m-1}|}$
$>|\alpha_{m}|+\frac{a^{2}-a-1}{a}-\frac{1}{a}-\dots-\frac{1}{a}$
$=|\alpha_{m}|+\frac{a^{2}-a-1}{a}-\frac{m-1}{a}=|\alpha_{m}|+\frac{a^{2}-a-m}{a}.$
It is easy to verify that $a^{2}-a-m=0$ and therefore, $q(R_{1})>|\alpha_{m}|>0$. The intermediate value theorem ensures that $q(x)$ has a root in $(|\alpha_{m}|,R_{1})$. But $R_{2}$ is the only root of $q(x)$ such that $|\alpha_{m}|<R_{2}$. Therefore, $R_{2}\in(|\alpha_{m}|,R_{1})$. Hence, $R_{2}<R_{1}$. ∎
Theorem 3.2.
Let $T(\lambda)=-B_{0}+I\lambda+\frac{B_{1}}{\lambda-\alpha_{1}}+\dots+\frac{B_{m}}{\lambda-\alpha_{m}}$, where $B_{i}$’s are $n\times n$ unitary matrices and $\alpha_{i}$’s are distinct complex numbers. Then the bound given in Theorem 2.10 is better than the bound given in Theorem 2.3.
Proof.
Let $R_{1}=a+|\alpha_{m}|$ be the bound for the eigenvalues of $T(\lambda)$ given in Theorem 2.3, where $a=\displaystyle\Bigg{[}\frac{(2m+1)+(4m+1)^{1/2}}{2}\Bigg{]}^{1/2}$. Let $R_{4}$ be the bound on the eigenvalues of $T(\lambda)$ obtained in Theorem 2.10. Since $B_{i}$’s are unitary matrices, $R_{4}=\displaystyle\frac{1}{2}\Big{(}1+|\alpha_{m}|+\sqrt{(|\alpha_{m}|-1)^{2}+4m}\Big{)}$. If $m=1$ and $\alpha_{1}=0$, then it is easy to check that $R_{1}=R_{4}$. Consider $m\geq 1$ and $\alpha_{m}\neq 0$. Define a scalar rational function $w(x):=x-1-\displaystyle\frac{m}{x-|\alpha_{m}|}$. The only zeros of $w(x)$ are $R_{4}^{\prime}=\displaystyle\frac{1}{2}\Big{(}1+|\alpha_{m}|-\sqrt{(|\alpha_{m}|-1)^{2}+4m}\Big{)}$ and $R_{4}=\displaystyle\frac{1}{2}\Big{(}1+|\alpha_{m}|+\sqrt{(|\alpha_{m}|-1)^{2}+4m}\Big{)}$. By Lemma 2.6, $R_{4}^{\prime}\in(-\infty,|\alpha_{m}|)$ and $R_{4}\in(|\alpha_{m}|,\infty)$. The remaining part of the proof follows as in the previous theorem. ∎
Some remarks are in order. In what follows, we work with the spectral norm.
Remark 3.3.
  1. (1)
    In Theorems 3.1 and 3.2, if the matrices are not unitary, we cannot determine which theorem gives the better bound. For example, consider $T(\lambda)=\displaystyle-B_{0}+I\lambda+\frac{B_{1}}{\lambda-\alpha_{1}}$, where $B_{0}=\begin{bmatrix}0&0\\ 1&0\end{bmatrix}$ and $B_{1}=\begin{bmatrix}1&0\\ 0&0\end{bmatrix}$.
    • a.
      If $\alpha_{1}=0.1$, then the bound given in Theorem 2.2 is $R_{1}=1.51$ and the bound given in Theorem 2.4 is $R_{2}=1.65$. Therefore, $R_{1}<R_{2}$.
    • b.
      If $\alpha_{1}=1$, then the bound given in Theorem 2.2 is $R_{1}=2.41$ and the bound given in Theorem 2.4 is $R_{2}=2$. Therefore, $R_{2}<R_{1}$.
    • c.
      Again, if $\alpha_{1}=0.1$, the bound given in Theorem 2.10 is $R_{4}=1.65$. Therefore, $R_{1}<R_{4}$. If $\alpha=1$ the bound given in Theorem 2.10 is $R_{4}=2$. Hence, $R_{4}<R_{1}$.
  2. (2)
    We cannot say which theorem gives a better bound between Theorem 2.3 and Theorem 2.8, even when the coefficients of $T(\lambda)$ are unitary matrices. For example consider, $T(\lambda)=\displaystyle-B_{0}+I\lambda+\frac{B_{1}}{\lambda-\alpha_{1}}$, where $B_{0}=B_{1}=I_{2}$.
    • a.
      If $\alpha_{1}=1$, then the bound given in Theorem 2.3 is $R_{1}=2.62$ and the bound given in Theorem 2.8 is $R_{3}=1+\sqrt{3}=2.73$. Therefore, we have $R_{1}<R_{3}$.
    • b.
      If and $\alpha_{1}=i$, then the bound given in Theorem 2.3 is $R_{1}=2.62$ and the bound given in Theorem 2.8 is $R_{3}=2.09$. In this case, $R_{3}<R_{1}$.
  3. (3)
    The same phenomenon happens with Theorems 2.4 and 2.8. Consider the same example as in $(2)$.
    • a.
      If $\alpha_{1}=-1.5$, then the bounds given in Theorems 2.4 and 2.8 are $R_{2}=2.28$ and $R_{3}=1$ respectively. Thus, $R_{3}<R_{2}$.
    • b.
      If $\alpha_{1}=1.5$, then the bounds given in Theorems 2.4 and 2.8 are $R_{2}=2.28$ and $R_{3}=3.27$ respectively. In this case, we have $R_{2}<R_{3}$.
  4. (4)
    The bounds obtained in Theorem 2.10 and Theorem 2.8 are also not comparable. Consider the same example given in $(2)$.
    • a.
      If $\alpha_{1}=1$, then bounds obtained in Theorems 2.10 and 2.8 are $R_{4}=2$ and $R_{3}=2.73$ respectively. Therefore, $R_{4}<R_{3}$.
    • b.
      If $\alpha_{1}=-0.5$, then bounds obtained in Theorems 2.10 and 2.8 are $R_{4}=1.78$ and $R_{3}=1$ respectively. Therefore, $R_{3}<R_{4}$.
4. Numerical results
Bounds on the moduli of eigenvalues of rational matrices are less studied in the literature. Recently in [4], the authors discuss some interesting techniques to derive eigenvalue bounds for general rational matrices. There are methods to determine approximate eigenvalues in specific regions using iterative methods and rational approximation methods (see for instance, [19], [20]), which are entirely different problems from ours. However, one can convert the rational matrix to a matrix polynomial and use existing results in the literature on matrix polynomials to compare these bounds. We do this and compare our bounds with the bounds given in [16]. We present three examples of REPs and compare the bounds obtained in Section 2 with the bounds given in [4] and [16]. In Example 4.1, we see that some of the bounds obtained in this manuscript are better than the bounds given in [4] and [16]. However, in general, one cannot determine which method gives a better bound.
Example 4.1.
Let $T(\lambda)=-B_{0}+I\lambda+\displaystyle\frac{B_{1}}{\lambda-0.1}$, where $I$ is the identity matrix of size $3$, $B_{0}=\begin{bmatrix}2&-1&0\\ -1&2&-1\\ 0&-1&1\end{bmatrix}$ and $B_{1}=\begin{bmatrix}-1&0&1\\ 0&-1&1\\ -1&0&1\end{bmatrix}$. Note that the maximum of the moduli of eigenvalues of $T(\lambda)$ is $3.54$.
Results Bounds Results Bounds
Theorem 2.2 3.70 Theorem $3.2$ of [16] 4.15
Theorem 2.4 (Theorem $3.8$, [4]) 3.83 Corollary $3.2.1$ of [16] 5.32
Theorem 2.8 3.90 Theorem $3.3$ of [16] 4.35
Theorem 2.11 4.36 Theorem $3.4$ of [16] 4.12
Theorem $3.9(1)$ of [4] 5.42 Corollary $3.4.2$ of [16] 5.07
Theorem $3.9(2)$ of [4] 4.25 Corollary $3.4.4$ of [16] 3.99
Theorem $3.9(3)$ of [4] 3.91 Corollary $3.4.6$ of [16] 4.91
Corollary $3.11$ of [4] 5.37 Theorem $3.6$ of [16] 4.35
Table 1. Bounds obtained from Section 2 and references [4], [16].
From Table 1, we can conclude that the bound obtained using Theorem 2.2 in this manuscript is better than other bounds.
Example 4.2.
Let $T(\lambda)=-B_{0}+I\lambda+\displaystyle\frac{B_{1}}{\lambda-2}$, where $I$ is the identity matrix of size $3$, $B_{0}=\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&2\end{bmatrix}$ and $B_{1}=\begin{bmatrix}0&0&0\\ 0&0&0\\ 0&0&-1\end{bmatrix}$. The maximum of the moduli of eigenvalues of $T(\lambda)$ is $3.00$.
Results Bounds Results Bounds
Theorem 2.2 4.41 Theorem $3.2$ of [16] 4.83
Theorem 2.4 (Theorem $3.8$, [4]) 3.00 Corollary $3.2.1$ of [16] 7.99
Theorem 2.8 4.65 Theorem $3.3$ of [16] 5.00
Theorem 2.11 3.00 Theorem $3.4$ of [16] 4.79
Theorem $3.9(1)$ of [4] 3.00 Corollary $3.4.2$ of [16] 6.32
Theorem $3.9(2)$ of [4] 3.00 Corollary $3.4.4$ of [16] 4.14
Theorem $3.9(3)$ of [4] 3.00 Corollary $3.4.6$ of [16] 5.12
Corollary $3.11$ of [4] 3.50 Theorem $3.6$ of [16] 5.00
Table 2. Bounds obtained from Section 2 and references [4], [16].
In Example 4.2, the bound obtained from Theorem 2.11 of this manuscript is better and it actually coincides with the bounds obtained from Theorems $3.8$ and $3.9$ of [4].
Example 4.3.
The following example arises in the finite element discretization of a boundary problem describing the eigenvibration of a string with a load of mass attached by an elastic spring. We refer readers to [5] for details about this particular REP. Let $T(\lambda)v=\left(A-B\lambda+C\frac{\lambda}{\lambda-\alpha}\right)v,$ where $A=\begin{bmatrix}6&-3&0\\ -3&6&-3\\ 0&-3&3\end{bmatrix}$, $B=\frac{1}{18}\begin{bmatrix}4&1&0\\ 1&4&1\\ 0&1&2\end{bmatrix}$, $C=\begin{bmatrix}0&0&0\\ 0&0&0\\ 0&0&1\end{bmatrix}$ and $\alpha=1$. Note that $B$ is invertible, therefore the above REP is the same as $\left(-(A+C)B^{-1}+I\lambda-CB^{-1}\frac{\alpha}{\lambda-\alpha}\right)v=0$. The maximum of the moduli of eigenvalues of $T(\lambda)$ is $94.60$.
Results Bounds Results Bounds
Theorem 2.2 98.46 Theorem $3.2$ of [16] 99.24
Theorem 2.4 (Theorem $3.8$, [4]) 97.38 Corollary $3.2.1$ of [16] 191.47
Theorem 2.8 99.18 Theorem $3.3$ of [16] 99.25
Theorem 2.11 98.46 Theorem $3.4$ of [16] 98.47
Theorem $3.9(1)$ of [4] 108.04 Corollary $3.4.2$ of [16] 101.13
Theorem $3.9(2)$ of [4] 98.27 Corollary $3.4.4$ of [16] 97.33
Theorem $3.9(3)$ of [4] 97.63 Corollary $3.4.6$ of [16] 98.33
Corollary $3.11$ of [4] 146.40 Theorem $3.6$ of [16] 99.25
Table 3. Bounds obtained from Section 2 and references [4], [16].
Table 3 suggests that the bound obtained from Corollary $3.4.4$ of [16] is sharper than other bounds.
In all three examples the upper bound obtained using Theorem $2.1(4)$ of [6] coincides with the bound given in Theorem 2.8. Let us point out that Roy and Bora [18] also study the eigenvalue location of quadratic matrix polynomials and compare their bounds with that of [6]; however, the bounds in [6] are better than that of [18]. In order to find a lower bound using Theorem 2.7, the coefficient matrices should satisfy the hypothesis given in the theorem. Note that the coefficient matrices of rational matrices given in Examples 4.1 and 4.3 do not satisfy the hypothesis of Theorem 2.7. Therefore, we find a lower bound only for Example 4.2. Note that the minimum of the moduli of eigenvalues of $T(\lambda)$ given in Example 4.2 is $1$. A lower bound on the moduli of eigenvalues obtained from Theorem 2.7 is $0.38$.
Data availability: No data was used for the research done in this article.
Declarations: The authors declare that there is no conflict of interest in this work.
Acknowledgements: Pallavi Basavaraju and Shrinath Hadimani are grateful for the financial assistance provided by the Council of Scientific and Industrial Research (CSIR) and the University Grants Commission (UGC), both departments of the Indian government.